### v1

# MSTN: Fast and Efficient Multivariate Time Series Model

Abstract
Real-world time-series data is highly non stationary and complex in dynamics that operate across multiple timescales, ranging from fast, short-term changes to slow, long-term trends. Most existing models rely on fixed-scale structural priors, such as patch-based tokenization, fixed frequency transformations, or frozen backbone architectures. This often leads to over-regularization of temporal dynamics, which limits their ability to adaptively model the full spectrum of temporal variations and impairs their performance on unpredictable, Sudden, high-magnitude events. To address this, we introduce the Multi-scale Temporal Network (MSTN), a novel deep learning architecture founded on a hierarchical multi-scale and sequence modeling principle. The MSTN framework integrates three core components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid for local patterns (ii) a sequence modeling component for long-range temporal dependencies. We empirically validate this with Bidirectional Long Short-Term Memory (LSTM) and Transformer variants, establishing a flexible foundation for future architectural advancements. and (iii) a gated fusion mechanism augmented with squeeze-and-excitation (SE) and multi-head temporal attention (MHTA) for dynamic, context-aware feature integration. This design enables MSTN to adaptively model temporal patterns from milliseconds to long-range dependencies within a unified framework. Extensive evaluations across time-series long-horizon forecasting, imputation, classification and generalizability study demonstrate that MSTN achieves competitive state-of-the-art (SOTA) performance, showing improvements over contemporary approaches including EMTSF, LLM4TS, HiMTM, TIME-LLM, MTST, SOFTS, iTransformer, TimesNet, and PatchTST. In total, MSTN establishes new SOTA performance on 24 of 32 benchmark datasets, demonstrating its consistent performance across diverse temporal tasks.
keywords:
Multivariate time series , multi-scale temporal modeling , time series forecasting , classification , imputation , computational efficiency , deep learning architectures
[…]
[inst1]organization=Department of Computer Science and Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
[…]
[inst2]organization=Department of Civil Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
1 Introduction
Multivariate time series analysis is a crucial component of modern AI, with significant real-world applications ranging from healthcare monitoring and industrial predictive maintenance to behavioral analytics. It has been extensively studied for many years and has been widely used in tasks such as weather forecasting, imputing missing data for data mining, and classification of trajectories for action recognition [wu2022autoformer, zhao2024himtmhierarchicalmultiscalemasked]. Due to its significant practical value, time series analysis has attracted considerable attention. Unlike language or video, time-series data consists of individual scalar measurements, where single time points rarely carry sufficient semantic meaning. This fundamental characteristic forces models to extract information from temporal variation patterns of continuity, periodicity, and trend, making robust temporal modeling exceptionally challenging [wu2023timesnettemporal2dvariationmodeling, lim2021time].
Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving SOTA results [nie2023timeseriesworth64]. Recent findings show that simple linear models can outperform sophisticated Transformer variants on common benchmarks [zeng2022transformerseffectivetimeseries], highlighting fundamental limitations in contemporary architectures. This is particularly critical in safety-sensitive domains like behavioral analytics and industrial monitoring, where patterns are often non-periodic and evolve across multiple timescales. The design of temporal models faces the constant challenge of balancing model strength with computational efficiency. Models like recurrent neural networks (RNN) and their gated variants [6795963, lai2018modelinglongshorttermtemporal] offer a strong sequential modeling but suffer from vanishing gradients and limited parallelism. Temporal convolutional networks (TCN)  [he2019temporal] improve computational efficiency but are constrained by fixed receptive fields. The Transformer architecture  [vaswani2023attentionneed] originally promised global dependency modeling, but its quadratic complexity prompted a line of research focused primarily on sparsification. Models like Informer [zhou2021informer] and FEDformer [zhou2022fedformer] develop efficient attention variants, yet often retain a point-wise view of time series, treating each time step as an independent token and struggling to capture local semantic structures.
These challenges have shaped several predominant, yet ultimately constrained, modern paradigms. One approach, exemplified by PatchTST [nie2023timeseriesworth64], adapts the patching technique from Vision Transformers, grouping adjacent time points into sub-series tokens. This approach successfully captures local semantics and benefits from channel-independent processing, but its reliance on a fixed patch size creates a fundamental rigidity, locking the model into a single, pre-defined scale of analysis. Another paradigm, improved by TimesNet [wu2023timesnettemporal2dvariationmodeling], reframes the problem by transforming 1D time series into 2D tensors based on identified periods. While powerful for data with strong periodic components, this approach is intrinsically dependent on reliable periodicity detection and the 2D transformation prior, making it less suited for non-stationary, irregular, or non-periodic patterns where such clear structure is absent. Building on these concepts, MTST [zhang2024multiresolutiontimeseriestransformerlongterm] employs a multi-resolution patch-based design with a multi-branch architecture for simultaneous modeling of diverse temporal patterns at different resolutions. However, its fixed choice of resolutions is less flexible and may fail to generalize across datasets with irregular or heterogeneous temporal patterns. Patch tokenization can also lose fine-grained dynamics, while the reliance on relative positional encoding restricts robustness in non-seasonal or event-driven sequences.
We identify a fundamental limitation on architectures that impose specific single-scale structural constraints, such as a fixed patch size, a period-based 2D shape, or frozen backbone architectures. These prior models struggle with the inherently multi-scale nature of many real-world temporal incidents. Their architecture limits their ability to model discriminative patterns that manifest simultaneously across different timescales, from brief micro-events to slow macro and long-term trends, such as found in risky driving behaviour. This gap highlights the need for a more flexible temporal representation that can primarily adapt to this hierarchical structure without relying on a single, potentially misaligned, structural assumption.
To address these challenges, we propose a DL model (MSTN) that naturally captures temporal hierarchies without relying on fixed and repeated structural assumptions. MSTN unifies multi-scale feature learning with global sequence modeling through a dual-path architecture with a sequence modeling core. We empirically validate this design with two instantiations: i) MSTN–BiLSTM, which excels at modeling local continuity and behavioral smoothness, and ii) MSTN–Transformer, which captures local complex patterns and long-range contextual dependencies. In both variants, convolutional encoders extract fine-grained local features, while the temporal core captures long-range dependencies. Their outputs are integrated via a gated fusion mechanism enhanced with SE recalibration and MHTA, yielding compact, discriminative, and temporally coherent representations. This design provides an unprecedented trade-off between representational capacity and computational efficiency.
We demonstrate that MSTN establishes new SOTA performance on 24 of 32 benchmark datasets for time series analysis. This includes 9 forecasting benchmarks, and 6 imputation tasks and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications. MSTN establishes a new SOTA in long-term forecasting, attaining the lowest prediction error on 6 out of 9 standard benchmarks and consistently exceeds leading architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], GPT2(6) [zhou2023fitsallpowergeneraltime], SOFTS [han2024softsefficientmultivariatetime] iTransformer [liu2024itransformerinvertedtransformerseffective], and PatchTST [nie2023timeseriesworth64]. Complementing these results, the MSTN-Transformer variant achieves first-place rankings 31 out of 48 ($\approx$65%) in all 48 imputation scenarios across the ETT, Electricity, and Weather datasets, demonstrating strong resilience under up to 50% missingness. evaluation scenarios and secure the highest average score on 5 out of 6 datasets. In the classification task, MSTN achieves improved performance over recent SOTA architectures such as GTP2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], LigthtTS  [zhang2022morefastmultivariatetime] and FEDformer [zhou2022fedformer]. We further evaluate MSTN for generalizability across 7 benchmark datasets from various domains like human safety, physical activity recognition, animal-welfare monitoring, and mechanical prognostics. Without domain-specific tuning, surprisingly, MSTN achieves uniformly strong performance, ranks first across all datasets, establishing new benchmarks and illustrating its robustness under diverse temporal dynamics. Finally, MSTN-Transformer achieves unprecedented performance with 0.020 MSE, 2.87 MB footprint and 0.72 ms inference time, representing a 22.4× accuracy improvement and 34.7× speedup over SOFTS [han2024softsefficientmultivariatetime]. The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.155 ms) significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al. [DARUS33012023] dataset. While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN-Transformer as an ideal solution for real-time and edge-AI temporal modeling applications. Our comprehensive evaluation establishes MSTN as a new benchmark, achieving SOTA performance on 24 of 32 datasets across forecasting, imputation, and classification tasks.
2 Results
This section presents a comprehensive evaluation of the proposed MSTN. Demonstrating its SOTA performance in the long-term forecasting, imputation, classification, and generalizability study. MSTN (MSTN-Transformer and MSTN-BiLSTM) achieves competitive performance, showing improvements over recent DL, Transformer based, and time series approaches.
2.1 Long-Term Forecasting
Table 1: Multivariate forecasting results with prediction lengths $H\in\{96,192,336,720\}$ for others and for ILI $H\in\{24,36,48,60\}$. Red/Blue: First/Second ranks.
Models MSTN-Trans. MSTN BiL. EMTSF LLM4TS HiMTM TIME-LLM MTST SOFTS iTransf.
Ours Ours 2025 2025 2024 2024 2024 2024 2024
H-Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTm1
96 0.052 0.174 0.168 0.334 0.271 0.325 0.285 0.343 0.280 0.331 0.272 0.334 0.286 0.338 0.325 0.361 0.334 0.368
192 0.103 0.242 0.111 0.258 0.322 0.351 0.324 0.366 0.321 0.357 0.310 0.358 0.327 0.366 0.375 0.389 0.377 0.391
336 0.127 0.271 0.197 0.350 0.350 0.370 0.353 0.385 0.347 0.378 0.352 0.384 0.362 0.389 0.405 0.412 0.426 0.420
720 0.151 0.301 0.166 0.314 0.414 0.404 0.408 0.419 0.395 0.411 0.383 0.411 0.414 0.421 0.466 0.447 0.491 0.459
Avg 0.123 0.262 0.161 0.314 0.339 0.362 0.342 0.378 0.336 0.369 0.329 0.372 0.347 0.378 0.393 0.403 0.407 0.410
ETTm2
96 0.201 0.325 0.238 0.390 0.156 0.240 0.165 0.254 0.164 0.254 0.161 0.253 0.162 0.251 0.180 0.261 0.180 0.264
192 0.373 0.467 0.426 0.514 0.212 0.280 0.220 0.292 0.221 0.291 0.219 0.293 0.220 0.291 0.246 0.306 0.250 0.309
336 0.345 0.477 0.517 0.594 0.263 0.315 0.268 0.326 0.273 0.326 0.271 0.329 0.272 0.326 0.319 0.352 0.311 0.348
720 0.411 0.509 0.576 0.614 0.351 0.371 0.350 0.380 0.355 0.378 0.352 0.379 0.358 0.379 0.405 0.401 0.412 0.407
Avg 0.333 0.445 0.439 0.528 0.245 0.301 0.251 0.313 0.253 0.312 0.251 0.313 0.253 0.312 0.287 0.330 0.288 0.332
ETTh1
96 0.142 0.285 0.298 0.459 0.359 0.384 0.371 0.394 0.355 0.386 0.362 0.392 0.358 0.390 0.381 0.399 0.386 0.405
192 0.144 0.293 0.288 0.446 0.399 0.411 0.403 0.412 0.401 0.417 0.398 0.418 0.396 0.414 0.435 0.431 0.441 0.436
336 0.152 0.307 0.163 0.314 0.418 0.422 0.420 0.422 0.420 0.429 0.430 0.427 0.391 0.420 0.480 0.452 0.487 0.458
720 0.181 0.349 0.223 0.383 0.436 0.454 0.422 0.444 0.425 0.447 0.442 0.457 0.430 0.457 0.499 0.488 0.503 0.491
Avg 0.155 0.309 0.243 0.401 0.403 0.417 0.404 0.418 0.400 0.419 0.408 0.423 0.394 0.420 0.449 0.442 0.454 0.447
ETTh2
96 0.297 0.429 0.384 0.493 0.262 0.324 0.269 0.332 0.273 0.334 0.268 0.328 0.257 0.326 0.297 0.347 0.297 0.349
192 0.300 0.436 0.401 0.505 0.328 0.371 0.328 0.377 0.334 0.371 0.329 0.375 0.309 0.361 0.373 0.394 0.380 0.400
336 0.377 0.495 0.383 0.501 0.347 0.387 0.353 0.396 0.353 0.398 0.368 0.409 0.302 0.366 0.410 0.426 0.428 0.432
720 0.400 0.509 0.445 0.530 0.372 0.420 0.383 0.425 0.371 0.412 0.381 0.417 0.372 0.416 0.411 0.433 0.427 0.445
Avg 0.349 0.472 0.403 0.507 0.329 0.374 0.333 0.382 0.332 0.379 0.334 0.383 0.310 0.367 0.373 0.400 0.383 0.407
ECL
96 0.041 0.161 0.558 0.585 0.126 0.217 0.128 0.223 0.129 0.220 0.131 0.224 0.127 0.222 0.143 0.233 0.148 0.240
192 0.042 0.163 0.372 0.464 0.144 0.234 0.146 0.240 0.147 0.238 0.152 0.241 0.144 0.238 0.158 0.248 0.162 0.253
336 0.042 0.162 0.382 0.471 0.158 0.248 0.163 0.258 0.157 0.249 0.160 0.248 0.162 0.256 0.178 0.269 0.178 0.269
720 0.047 0.173 0.469 0.521 0.190 0.277 0.200 0.292 0.198 0.285 0.192 0.298 0.199 0.289 0.218 0.305 0.225 0.317
Avg 0.043 0.165 0.445 0.510 0.154 0.244 0.159 0.253 0.157 0.248 0.158 0.252 0.158 0.251 0.174 0.264 0.178 0.270
Traffic
96 0.017 0.110 0.085 0.202 0.343 0.225 0.372 0.259 0.358 0.240 0.362 0.248 0.356 0.244 0.376 0.251 0.395 0.268
192 0.018 0.111 0.086 0.205 0.369 0.238 0.391 0.265 0.368 0.248 0.374 0.247 0.375 0.251 0.398 0.261 0.417 0.276
336 0.020 0.112 0.086 0.204 0.382 0.242 0.405 0.275 0.379 0.250 0.385 0.271 0.386 0.256 0.415 0.269 0.433 0.283
720 0.020 0.113 0.087 0.207 0.424 0.270 0.437 0.292 0.430 0.276 0.430 0.288 0.425 0.279 0.447 0.287 0.467 0.302
Avg 0.019 0.111 0.086 0.205 0.379 0.243 0.401 0.273 0.384 0.254 0.388 0.264 0.386 0.257 0.409 0.267 0.428 0.282
Weather
96 0.109 0.263 0.110 0.264 0.138 0.177 0.147 0.196 0.141 0.182 0.147 0.201 0.150 0.199 0.166 0.208 0.174 0.214
192 0.111 0.264 0.111 0.265 0.181 0.220 0.191 0.238 0.188 0.228 0.189 0.234 0.194 0.240 0.217 0.253 0.221 0.254
336 0.114 0.268 0.115 0.269 0.230 0.260 0.241 0.277 0.240 0.273 0.262 0.279 0.246 0.281 0.282 0.300 0.278 0.296
720 0.128 0.286 0.132 0.289 0.304 0.315 0.313 0.329 0.312 0.322 0.304 0.316 0.319 0.333 0.356 0.351 0.358 0.347
Avg 0.115 0.270 0.117 0.272 0.213 0.243 0.223 0.260 0.220 0.251 0.225 0.257 0.227 0.263 0.255 0.278 0.258 0.278
Exchange
96 0.495 0.596 0.231 0.401 - - - - - - - - - - - - - -
192 0.456 0.566 0.278 0.437 - - - - - - - - - - - - - -
336 0.906 0.803 0.303 0.469 - - - - - - - - - - - - - -
720 0.815 0.734 0.818 0.795 - - - - - - - - - - - - - -
Avg 0.668 0.675 0.408 0.526 - - - - - - - - - - - - - -
ILI
24 0.129 0.254 0.161 0.328 1.617 0.732 - - - - 1.285 0.727 - - - - - -
36 0.155 0.287 0.167 0.333 1.586 0.728 - - - - 1.404 0.814 - - - - - -
48 0.180 0.323 0.168 0.334 1.587 0.753 - - - - 1.523 0.807 - - - - - -
60 0.190 0.332 0.169 0.334 1.560 0.768 - - - - 1.531 0.854 - - - - - -
Avg 0.165 0.302 0.166 0.332 1.588 0.745 - - - - 1.436 0.801 - - - - - -
We evaluate long-term forecasting performance using established protocols from the prior literature [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64] across nine diverse benchmark datasets. The prediction horizons are set to $H\in\{96,192,336,720\}$ for the standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI data set. As presented in Table 1, the MSTN-Transformer and MSTN BiLSTM shows improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm],SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
[…]
2.2 Imputation
Table 2: Imputation Task Results Comparison. We randomly mask 12.5%, 25%, 37.5%, and 50% time points to compare the model performance under different missing degrees. Red/Blue: First/Second ranks.
Models MSTN-Tra. MSTN BiL. GPT2(3) TimesNet PatchTST LightTS DLinear Stationary
(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
ETTh1
12.5% 0.024 0.040 0.032 0.047 0.043 0.140 0.057 0.159 0.093 0.201 0.240 0.345 0.151 0.267 0.060 0.165
25% 0.050 0.083 0.067 0.096 0.054 0.156 0.069 0.178 0.107 0.217 0.265 0.364 0.180 0.292 0.080 0.189
37.5% 0.083 0.131 0.103 0.147 0.072 0.180 0.084 0.196 0.120 0.230 0.296 0.382 0.215 0.318 0.102 0.212
50% 0.127 0.187 0.148 0.204 0.107 0.216 0.102 0.215 0.141 0.248 0.334 0.404 0.257 0.347 0.133 0.240
Avg 0.071 0.110 0.088 0.124 0.069 0.173 0.078 0.187 0.115 0.224 0.284 0.373 0.201 0.306 0.094 0.201
ETTh2
12.5% 0.009 0.024 0.022 0.040 0.039 0.125 0.040 0.130 0.057 0.152 0.101 0.231 0.100 0.216 0.042 0.133
25% 0.022 0.054 0.043 0.078 0.044 0.135 0.046 0.141 0.061 0.158 0.115 0.246 0.127 0.247 0.049 0.147
37.5% 0.038 0.089 0.069 0.121 0.051 0.147 0.052 0.151 0.067 0.166 0.126 0.257 0.158 0.276 0.056 0.158
50% 0.066 0.135 0.102 0.170 0.059 0.158 0.060 0.162 0.073 0.174 0.136 0.268 0.183 0.299 0.065 0.170
Avg 0.034 0.076 0.059 0.102 0.048 0.141 0.049 0.146 0.065 0.163 0.119 0.250 0.142 0.259 0.053 0.152
ETTm1
12.5% 0.024 0.038 0.036 0.049 0.017 0.085 0.019 0.092 0.041 0.130 0.075 0.180 0.058 0.162 0.026 0.107
25% 0.050 0.080 0.072 0.099 0.022 0.096 0.023 0.101 0.044 0.135 0.093 0.206 0.080 0.193 0.032 0.119
37.5% 0.082 0.124 0.112 0.150 0.029 0.111 0.029 0.111 0.049 0.143 0.113 0.231 0.103 0.219 0.039 0.131
50% 0.121 0.176 0.160 0.210 0.040 0.128 0.036 0.124 0.055 0.151 0.134 0.255 0.132 0.248 0.047 0.145
Avg 0.069 0.104 0.095 0.127 0.028 0.105 0.027 0.107 0.047 0.140 0.104 0.218 0.093 0.206 0.036 0.126
ETTm2
12.5% 0.039 0.048 0.052 0.057 0.017 0.076 0.018 0.080 0.026 0.094 0.034 0.127 0.062 0.166 0.021 0.088
25% 0.080 0.097 0.107 0.116 0.020 0.080 0.020 0.085 0.028 0.099 0.042 0.143 0.085 0.196 0.024 0.096
37.5% 0.132 0.153 0.164 0.176 0.022 0.087 0.023 0.091 0.030 0.104 0.051 0.159 0.106 0.222 0.027 0.103
50% 0.187 0.215 0.225 0.241 0.025 0.095 0.026 0.098 0.034 0.110 0.059 0.174 0.131 0.247 0.030 0.108
Avg 0.109 0.128 0.137 0.147 0.021 0.084 0.022 0.088 0.029 0.102 0.046 0.151 0.096 0.208 0.026 0.099
ECL
12.5% 0.002 0.014 0.034 0.048 0.080 0.194 0.085 0.202 0.055 0.160 0.102 0.229 0.092 0.214 0.093 0.210
25% 0.005 0.028 0.069 0.096 0.087 0.203 0.089 0.206 0.065 0.175 0.121 0.252 0.118 0.247 0.097 0.214
37.5% 0.008 0.044 0.108 0.147 0.094 0.211 0.094 0.213 0.076 0.189 0.141 0.273 0.144 0.276 0.102 0.220
50% 0.013 0.063 0.150 0.202 0.101 0.220 0.100 0.221 0.091 0.208 0.160 0.293 0.175 0.305 0.108 0.228
Avg 0.007 0.037 0.090 0.123 0.090 0.207 0.092 0.210 0.072 0.183 0.131 0.262 0.132 0.260 0.100 0.218
Weather
12.5% 0.003 0.016 0.026 0.036 0.026 0.049 0.025 0.045 0.029 0.049 0.047 0.101 0.039 0.084 0.027 0.051
25% 0.007 0.032 0.051 0.070 0.028 0.052 0.029 0.052 0.031 0.053 0.052 0.111 0.048 0.103 0.029 0.056
37.5% 0.011 0.049 0.078 0.109 0.033 0.060 0.031 0.057 0.035 0.058 0.058 0.121 0.057 0.117 0.033 0.062
50% 0.015 0.068 0.111 0.153 0.037 0.065 0.034 0.062 0.038 0.063 0.065 0.133 0.066 0.134 0.037 0.068
Avg 0.009 0.041 0.066 0.092 0.031 0.056 0.030 0.054 0.033 0.056 0.055 0.117 0.052 0.110 0.032 0.059
As presented in Table 2, MSTN-Transformer achieves SOTA performance in 32 out of 48 ($\approx$66%) imputation evaluation scenarios and delivers the improved average results on 5 of the 6 benchmark datasets, while MSTN-BiLSTM demonstrates competitive performance. Following the TimesNet benchmark protocol [wu2023timesnettemporal2dvariationmodeling], we evaluate imputation performance on electricity and weather domain datasets, including ETT [zhou2021informer], ECL [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather]. To simulate real-world scenarios, we employ random masking ratios of 12.5%, 25%, 37.5%, 50%. This shows improved performance over several recent approaches, including GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
2.3 Time-Series Classification
Table 3 presents the classification results across 10 UEA standard benchmark datasets. The proposed MSTN framework is compared against SOTA time-series classification methods, including GPT-2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime]. MSTN demonstrates competitive performance in multiple datasets. MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
Table 3: Comprehensive evaluation of MSTN across diverse time-series benchmarks.
(a) Classification accuracy (%) comparison between MSTN (MSTN- Transformer and MSTN-BiLSTM) and SOTA baselines. Red/Blue: First/Second accuracy ranks.
Dataset MSTN MSTN GPT2 Times- Light DLinear Flow. ETS. FED. Stat. Auto. Py. In. Re. Trans.
Trans. BiL. (6) (Net) TS (2023) (2022) (2022) (2022a) (2022) (2021) (2021a) (2021) (2020) (2017)
(ours) (ours) (2023) (2023) (2022)
Ethanol 33.92 32.43 31.9 35.7 29.7 32.6 33.8 28.1 31.2 32.7 31.6 30.8 31.6 31.9 32.7
FaceDetect 63.20 60.30 67.3 68.6 67.5 68.0 67.6 66.3 66.0 68.0 68.4 65.7 67.0 68.6 67.3
Handwriting 43.0 35.50 32.0 32.1 26.1 27.0 33.8 32.5 28.0 31.6 36.7 29.4 32.8 27.4 32.0
Heartbeat 78.95 81.49 76.1 78.0 75.1 75.1 77.6 71.2 73.7 73.7 74.6 75.6 80.5 77.1 76.1
JapaneseV 100 100 98.6 98.4 96.2 96.2 98.9 95.9 98.4 99.2 96.2 98.4 98.9 97.8 98.7
PEMS-SF 94.32 96.59 82.1 89.6 88.4 75.1 83.8 86.0 80.9 87.3 82.7 83.2 81.5 82.7 82.1
SCP1 93.79 90.27 93.2 91.8 89.8 87.3 92.5 89.6 88.7 89.4 84.0 88.1 90.1 90.4 92.2
SCP2 62.42 60.53 59.4 57.2 51.1 50.5 56.1 55.0 54.4 57.2 50.6 53.3 53.3 56.7 53.9
SpokenArabic 99.32 99.15 99.2 99.0 100 81.4 98.8 100 100 100 100 99.6 100 97.0 98.4
UWave 76.00 68.18 88.1 85.3 80.3 82.1 86.6 85.0 85.3 87.5 85.9 83.4 85.6 85.6 85.6
(b) Generalizability Study: Performance comparison of MSTN-BiLSTM and MSTN-Transformer with TimesNet and PatchTST across seven international cross-domain datasets. Red/Blue: First/Second accuracy ranks; Cyan/Violet: First/Second inference time ranks.
MSTN TimesNet PatchTST
Dataset Det. Dom. Cnt. A.Block Acc.(%) F1 Prec. Rec. Size I.Time Acc.(%) I.Time Acc.(%) I.Time Prior Acc.
Rodeg.[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
Transf. 99.53 0.947 0.951 0.942 0.54 0.155
Boubez.[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
Transf. 91.76 0.927 0.921 0.934 0.96 0.24
UCI-HAR[har240] Act. Hum. IT. BiLSTM 96.59 0.965 0.966 0.965 7.14 2.03 91.38 45.30 93.21 4.36 83.35[Km,NB][Ismi2016KmeansCB]
Transf. 95.58 0.956 0.958 0.955 29.98 3.48
PAMAP2[pamap2_physical_activity_monitoring_231] Phy. Hum. US. BiLSTM 99.69 0.996 0.996 0.996 0.76 0.086 95.13 0.46 98.02 3.21 90.00[kNN][6246152]
Transf. 99.52 0.995 0.995 0.995 1.58 0.20
ActBeC.[DISSANAYAKE2025111462] Calf. Anim. IE. BiLSTM 93.00 0.923 0.923 0.923 4.14 1.76 90.65 8.53 62.45 2.20 84.00[RCCV][DISSANAYAKE2025111462]
Transf. 88.14 0.883 0.873 0.913 0.89 0.14
MetroPT3[metropt3] Met. Mech. PT. BiLSTM 93.10 0.930 0.930 0.931 0.57 0.042 93.00 0.09 81.67 0.19 62.00[SAE][9564181]
Transf. 93.83 0.938 0.940 0.938 1.30 0.19
NASA[NasaTurboFan] Eng. Mech. US. BiLSTM 20.24 4.25 1.48 15.56 4.52 31.59 2.31
Transf. 11.26 1.31 0.45
Abbreviations: Det: Details (Sim: Simulator PTW data, Real: Real PTW Data, Act: Activity human, Phy: Physical activity, Calf: Calf Behavior, Met: Metro predictive maintenance, Eng: Engine sensor), Dom: Domain (Hum: Human Safety, Anim: Animal welfare, Mech: Mechanical), Cnt: Country (IN: India, FR: France, DE: Germany, IT: Italy, US: USA, IE: Ireland, PT: Portugal), A.Block: Architecture block (Transf. Transformer), I.Time: Inference time, †RMSE values.
2.4 Generalizability Study
As summarized in Table 3, the proposed MSTN framework demonstrates remarkable generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics. Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64] in both accuracy and computational efficiency.
MSTN achieves competitive performance across fundamentally different domains, demonstrating its capabilities as a universal temporal model. In human safety applications, MSTN-Transformer reaches 99.53% accuracy on German collision prediction data[DARUS33012023], substantially outperforming prior work (91.0% RF,GB [Rodegast2024]). For activity recognition, MSTN-BiLSTM achieves 96.59% accuracy on the UCI-HAR dataset[har240], representing a 13.2% improvement over TimesNet. The framework also excels in real-world fall detection with 93.75% accuracy, surpassing dedicated DT-based methods. MSTN-BiLSTM also shows strong performance in animal welfare[DISSANAYAKE2025111462] assessment with 93.00% accuracy on calf behavior recognition. Furthermore, MSTN-Transformer demonstrates exceptional capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan dataset[NasaTurboFan] and significantly outperforming all baseline predictors in remaining useful life estimation.
MSTN delivers this performance with good computational efficiency. For instance, on the Rodegast [DARUS33012023] Germany data set, the MSTN-Transformer variant achieves 0.155 ms/sample inference latency 33× faster than TimesNet and 31.0× faster than PatchTST while maintaining compact model size. This efficiency enables deployment on resource-constrained edge devices for real-time safety applications. MSTN achieves consistent SOTA performance across fundamentally different domains from behavioral analytics to mechanical systems, without any architectural modifications.
2.5 Ablation Study
Table 4 presents an ablation study evaluating the effectiveness of the proposed MSTN-Transformer components across six datasets (ECL, Weather, ETTm1, ETTh2, Traffic, and ILI) and the MSTN-BiLSTM architecture components on the Weather, ETTm2, and Traffic datasets. To verify the contribution of each component, we compare the full MSTN-Transformer model against five variants: w/o CNN, which removes the convolutional pathway; w/o Transformer, which eliminates the core Transformer module; w/o SE, which removes the squeeze-and-excitation blocks; w/o MHTA, which excludes the Multi-Head Temporal Attention; and w/o Gated F., which removes the gated fusion mechanism. In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer to assess its importance in temporal sequence modeling. These results validate our architectural design choices and demonstrate the complementary nature of the integrated components.
Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$.
Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHTA w/o Gated F.
Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
MSTN Transformer
ECL
96 0.041 0.161 0.045 0.165 0.048 0.173 0.042 0.163 0.043 0.164 0.045 0.168
192 0.043 0.165 0.043 0.169 0.048 0.173 0.043 0.164 0.044 0.166 0.047 0.170
336 0.042 0.162 0.046 0.168 0.050 0.178 0.046 0.167 0.045 0.169 0.049 0.168
720 0.047 0.173 0.051 0.179 0.055 0.186 0.049 0.172 0.050 0.178 0.053 0.179
Weather
96 0.109 0.263 0.111 0.265 0.118 0.273 0.109 0.263 0.110 0.264 0.110 0.264
192 0.111 0.264 0.115 0.269 0.120 0.275 0.111 0.264 0.111 0.265 0.112 0.265
336 0.114 0.268 0.120 0.274 0.123 0.279 0.113 0.268 0.114 0.269 0.114 0.269
720 0.128 0.286 0.138 0.296 0.136 0.294 0.130 0.287 0.131 0.289 0.130 0.287
ETTm1
96 0.052 0.174 0.055 0.175 0.089 0.227 0.061 0.185 0.058 0.178 0.055 0.175
192 0.103 0.242 0.105 0.243 0.105 0.243 0.109 0.246 0.098 0.233 0.096 0.229
336 0.127 0.271 0.129 0.272 0.156 0.307 0.140 0.283 0.151 0.296 0.147 0.290
720 0.151 0.301 0.155 0.306 0.172 0.319 0.142 0.288 0.159 0.308 0.158 0.307
ETTh2
96 0.297 0.429 0.370 0.492 0.365 0.485 0.374 0.482 0.377 0.491 0.390 0.485
192 0.300 0.436 0.404 0.511 0.391 0.505 0.445 0.529 0.382 0.498 0.395 0.501
336 0.377 0.495 0.415 0.520 0.388 0.501 0.396 0.504 0.387 0.498 0.396 0.505
720 0.400 0.509 0.427 0.522 0.384 0.498 0.426 0.523 0.426 0.531 0.439 0.531
Traffic
96 0.017 0.110 0.020 0.113 0.026 0.128 0.019 0.112 0.021 0.116 0.020 0.112
192 0.018 0.111 0.020 0.115 0.026 0.127 0.020 0.113 0.023 0.116 0.021 0.116
336 0.020 0.112 0.022 0.118 0.029 0.135 0.021 0.115 0.024 0.118 0.023 0.114
720 0.020 0.113 0.026 0.128 0.026 0.128 0.022 0.118 0.025 0.119 0.025 0.120
ILI
24 0.129 0.254 0.140 0.272 0.139 0.274 0.132 0.256 0.131 0.253 0.133 0.256
36 0.155 0.287 0.167 0.301 0.168 0.307 0.161 0.298 0.157 0.293 0.168 0.304
48 0.180 0.323 0.208 0.344 0.193 0.332 0.187 0.327 0.183 0.325 0.191 0.328
60 0.190 0.332 0.197 0.338 0.199 0.339 0.193 0.333 0.193 0.332 0.194 0.338
MSTN BiLSTM
Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHTA w/o Gated F.
Weather
96 0.110 0.264 0.112 0.266 0.119 0.274 0.110 0.264 0.110 0.263 0.110 0.264
192 0.111 0.265 0.1132 0.267 0.120 0.275 0.112 0.265 0.111 0.265 0.112 0.266
336 0.115 0.269 0.116 0.270 0.123 0.279 0.115 0.269 0.115 0.269 0.115 0.269
720 0.132 0.289 0.138 0.292 0.139 0.296 0.132 0.293 0.133 0.295 0.134 0.298
ETTm2
96 0.238 0.390 0.243 0.400 0.264 0.430 0.262 0.394 0.240 0.388 0.261 0.399
192 0.426 0.514 0.501 0.560 0.528 0.617 0.436 0.540 0.419 0.535 0.411 0.512
336 0.517 0.594 0.581 0.623 0.571 0.612 0.535 0.592 0.536 0.593 0.626 0.655
720 0.576 0.614 0.592 0.635 0.578 0.619 0.681 0.661 0.829 0.738 0.641 0.647
Traffic
96 0.085 0.202 0.086 0.202 0.102 0.222 0.085 0.202 0.087 0.204 0.087 0.205
192 0.087 0.205 0.088 0.207 0.101 0.221 0.087 0.205 0.088 0.208 0.088 0.209
336 0.086 0.204 0.086 0.205 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
720 0.087 0.207 0.089 0.210 0.102 0.224 0.092 0.218 0.091 0.217 0.093 0.218
2.6 Model Complexity and Edge-AI Deployability
The MSTN-Transformer and MSTN-BiLSTM architectures exemplify a fundamental design trade-off in temporal modeling. Both architectures integrate efficient parallel components CNN pathways ($\mathcal{O}(T)$) with either Transformer ($\mathcal{O}(T^{2})$) or BiLSTM ($\mathcal{O}(T)$) mechanisms, combined with $\mathcal{O}(T)$ fusion operations and finalized by a multi-head temporal attention layer of $\mathcal{O}(T^{2})$ complexity. This quadratic component dictates the overall asymptotic complexity of $\mathcal{O}(T^{2})$. However, this design preserves robust temporal modeling capabilities and achieves improved performance and practical efficiency compared to pure $\mathcal{O}(T^{2})$ Transformer-based models. Linear models such as DLinear and TS-Mixer achieve optimal $\mathcal{O}(T)$ complexity, while spectral and frequency-based approaches such as FEDformer and TimesNet attain $\mathcal{O}(T)$ and $\mathcal{O}(T\log T)$ complexity, respectively. The recently proposed MTST further reduces the cost to $\mathcal{O}(J_{b_{n}^{\prime}}^{2})$ via patching. Although efficient, these methods often trade off capacity for unconstrained long-range dependency modeling.
Following the results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our results with their metrics reported in Table 5. The MSTN-Transformer achieves improved performance with 0.020 MSE and 0.72 ms inference time, representing 22.4× accuracy improvement and 34.7× speedup over SOFTS, achieving a favorable balance of accuracy and speed. This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
Table 5: A: Performance Analysis (Traffic Dataset, $L=96$, $H=720$). B: MSTN-Trans. Inference Time Component Contributions. Red/Blue: First/Second ranks.
MSE Perf. Gap Inf. Time (ms) Speed Gap
A: SOTA MSE and Efficiency
[…]
B: MSTN-Trans. Component Contributions Ablation Study (Traffic Dataset, $H=720$)
Variant Size (MB) Time (ms) Speed Gap Role
[…]
As shown in Table 3, MSTN-Transformer delivers competitive performance across domains, attaining 99.53% accuracy on Rodegast [DARUS33012023] while maintaining efficient inference at 0.155 ms per sample with only 0.54 MB model size. This represents up to 33× speedup over TimesNet (5.07 ms) and 31.0× over PatchTST (4.80 ms), demonstrating favorable performance-efficiency trade-offs across diverse datasets. The combination of compact model sizes (0.54–7.14 MB), improved accuracy, and millisecond-level latency positions MSTN as an ideal solution for complex multivariate time series analysis tasks in edge computing environments. These results demonstrate that capable temporal models can achieve practical deployability in edge environments. The small model footprint satisfies memory constraints, millisecond-level latency enables real-time operation, and high accuracy ensures reliability. This collectively addresses core edge-AI challenges for applications like safety monitoring, predictive maintenance, and activity recognition.
3 Discussion
This work introduces MSTN, a DL architecture founded on a hierarchical multi-scale modeling principle. It integrates three core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; sequential modeling component for global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a gated fusion mechanism augmented with SE and MHTA to enable dynamic, context-aware feature integration. Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures), forcing a compromise between scale diversity and inference speed. In contrast, our approach efficiently integrates convolutional local pattern extraction with global dependency modeling via an adaptive gated fusion mechanism. This enables our model to capture multi-scale temporal features without compromising computational efficiency, ultimately achieving state-of-the-art performance.
Comprehensive evaluations demonstrate that the MSTN-Transformer establishes a new SOTA across forecasting, imputation, classification, and generalizability. It achieves a 22.4× improvement in accuracy over the SOFTS model while being 34.7× faster in inference, all within a compact 2.87 MB footprint, while MSTN BiLSTM shows competitive performance. This dual advancement in both performance and efficiency is a key differentiator from prior work. The model also exhibits exceptional cross-domain generalization, with its small model size (0.54–7.14 MB) making it suitable for edge deployment. Ablation studies confirm that each component is critical, with the transformer core providing a foundational capability and the multi-scale and fusion modules delivering significant competitive gains.
Future work will explore cross-variate attention mechanisms, and large-scale pre-training to further enhance MSTN’s capabilities as a foundation model for time series understanding. Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time forecasting applications across diverse domains from healthcare to industrial monitoring.
4 Methods
4.1 Proposed Architecture: MSTN
4.1.1 Architecture Overview
(a) MSTN architecture
(b) MSTN multi-scale signal processing pipeline
Fig. 1: Proposed MSTN: (a) architectural diagram and (b) signal processing pipeline
The proposed MSTN (Fig. 1a) is a novel hybrid DL architecture founded on a hierarchical multi-scale modeling principle. It uses a dual-branch processing pipeline: (i) a convolutional branch that extracts detailed local motion patterns, and (ii) sequence modeling core that learns temporal dependencies in the time-series data, which we empirically validate using both BiLSTM and Transformer variant in forecasting, imputation, classification and generalizability studies. (iii) The outputs of both branches are integrated using a learnable gated fusion mechanism, followed by channel-wise recalibration through SE blocks and MHTA, enabling the model to focus on critical, risk-indicating time steps. The fused representation is normalized, regularized with dropout, and linearly projected to produce probabilities. Training takes advantage of focal loss to mitigate class imbalance. Fig. 1b shows the processing pipeline of time series signals. CNN pathway captures local patterns, while BilSTM and Transformer captures long-range temporal dependencies in the time-series data, with gated fusion adaptively combining both feature streams.
4.1.2 Model Structure
As aforementioned, based on the multi-scale nature of time series, we propose MSTN with a parallel architecture to capture temporal patterns at different resolutions. For comprehensive temporal modeling, we design dual pathways within MSTN that process the time series through complementary mechanisms and simultaneously model both local and global variations through a parameter-efficient novel hybrid block. The MSTN architecture and signal processing pipeline are illustrated in Fig. 1.
We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times M}$ with lookback window sequence length $L$ and $M$ variates features, MSTN learns rich temporal representations for multiple tasks including forecasting, imputation and classification through a novel parallel architecture.
Forward Process. The input $X\in\mathbb{R}^{B\times T\times D}$ is processed through dual complementary pathways that operate simultaneously. The CNN pathway captures local temporal patterns while the Transformer pathway models global dependencies or BiLSTM models sequential intelligence, with their outputs fused through sophisticated gated mechanisms for enhanced representation learning. To capture multi-scale temporal variations in length-$T$ time series with $D$ variates organized as $X_{1D}\in\mathbb{R}^{B\times T\times D}$, the architecture processes inputs through two parallel pathways with convolutional pathway and sequence modeling (Transformer or BiLSTM pathway).
The convolutional pathway captures fine-grained, short-term temporal patterns through hierarchical feature learning. The input undergoes successive 1D convolutions to extract multi-scale features, expanding the feature dimension while preserving temporal resolution, followed by refinement of the representations. To aggregate these temporal features into a compact representation, we apply global average pooling across the entire sequence:
$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{Conv1D}_{7}(X_{1D}))\in\mathbb{R}^{B\times 128\times T}$ (1)
$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times T}$ (2)
$\mathbf{z}_{\text{cnn}}$ $=\text{GlobalAvgPool1D}(H_{\text{conv}}^{(2)})\in\mathbb{R}^{B\times 64}$ (3)
where each of the 64 feature channels is summarized by its average activation over time through temporal pooling. The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training.
We empirically validate two complementary sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM. Features from both pathways are elegantly concatenated, uniting the CNN’s local precision with either Transformer or BiLSTM, to capture long-range temporal dependencies. When using Transformer: The architecture captures long-range dependencies through a multi-layer transformer encoder. The input is processed through self-attention mechanisms where each of the T time steps maintains a 64-dimensional representation enriched with global contextual information. To consolidate these temporal representations, we apply sequence mean pooling across the time dimension:
$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times T\times 64};\quad\mathbf{z}_{\text{trans}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{trans},t}\in\mathbb{R}^{B\times 64}$ (4)
The transformer encoder uses 4 layers with 8 attention heads. When using BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times T\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
The BiLSTM employs 2 layers with 64 hidden units per direction, capturing both causal and anti-causal temporal relationships. The CNN pathway utilizes global average pooling to capture local pattern summaries, while the Transformer and BiLSTM pathways employ sequence mean pooling to preserve global contextual information. The parallel features undergo sophisticated fusion through a multi-stage enhancement process that intelligently combines the strengths of both pathways. For MSTN-Transformer:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans}}]\in\mathbb{R}^{128}$ (6)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{64}$. For MSTN-BiLSTM:
$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{bilstm}}]\in\mathbb{R}^{192}$ (7)
where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{bilstm}}\in\mathbb{R}^{128}$. The architecture then performs adaptive weighting through an intelligent gated fusion mechanism, illustrated in Fig. 1, that learns to dynamically balance feature contributions:
$\mathbf{z}_{\text{fused}}=\mathbf{z}_{\text{concat}}\odot\sigma(W_{g}\mathbf{z}_{\text{concat}}+b_{g})$ (8)
where for MSTN-Transformer: $W_{g}\in\mathbb{R}^{128\times 128}$, $b_{g}\in\mathbb{R}^{128}$ and for MSTN-BiLSTM: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$. This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal intelligence. The fused features are reshaped for sequence processing: $z_{\text{fused\_seq}}=z_{\text{fused}}\text{.unsqueeze}(1)\text{.repeat}(1,T,1)\in\mathbb{R}^{B\times T\times d}$ where $d=128$ for MSTN-Transformer and $d=192$ for MSTN-BiLSTM. Channel-wise attention then enhances features through separate SE blocks for each variant: For MSTN-Transformer: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{16\times 128}$, $W_{2}\in\mathbb{R}^{128\times 16}$ with reduction ratio 8. For MSTN-BiLSTM: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8. The SE mechanism amplifies informative channels while suppressing less useful ones. Temporal relationships are refined through multi-head attention that captures interdependencies across the enhanced features:
$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(T^{-1}\textstyle\sum_{t=1}^{T}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$
with 4 attention heads (dim=64). The attention output undergoes layer normalization for stabilization and dropout for regularization, producing refined features $\mathbf{z}_{\text{final}}$ for the final classification head. The parallel pathway design enables MSTN to achieve comprehensive temporal modeling where $\mathcal{F}_{\text{CNN}}$ captures local patterns and $\mathcal{F}_{\text{Transformer}}$ models global dependencies, with hierarchical fusion creating synergistic representations that surpass individual pathway capabilities. This architectural innovation establishes a new paradigm for multi-scale time series analysis.
4.1.3 Method-Level Innovation
Many recent time-series models including PatchTST [nie2023timeseriesworth64] (patch-based tokenization), TimesNet [wu2023timesnettemporal2dvariationmodeling] (fixed frequency transformations), and LLM-based approaches like LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] (frozen backbone architectures) often rely on rigid, fixed-scale structural priors. This architectural inflexibility forces a fundamental trade-off: models either compromise on capturing multi-scale temporal diversity or suffer from high inference latency, rendering them suboptimal for real-time, resource-constrained deployment.
The proposed MSTN addresses this core challenge through a novel multi-scale hierarchical architecture designed to capture both local patterns and long-range temporal dependencies without structural compromises. The architecture integrates three key components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; (ii) a sequence temporal modeling core, which we empirically validate with both BiLSTM and Transformer variants, establishing a flexible foundation for future architectural advancements; and (iii) a gated fusion mechanism augmented with SE and MHTA for dynamic, context-aware feature integration. This hierarchical design enables seamless adaptation to diverse temporal modeling paradigms while maintaining architectural consistency. Our comprehensive experiments demonstrate that these design choices allow MSTN to achieve improved performance compared to SOTA baselines while supporting low-latency inference critical for edge deployment.
4.2 Dataset and Baselines
4.2.1 Long Term Forecasting
The proposed MSTN model was evaluated on eight widely-used benchmark datasets: Weather, Traffic, Electricity, ILI, and four variants of ETT (ETTh1, ETTh2, ETTm1, ETTm2) [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64]. These datasets are standard in the field and have been extensively used in prior studies, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime] and iTransformer [liu2024itransformerinvertedtransformerseffective]. Notably, large-scale datasets such as Weather, Traffic, and Electricity contain a significantly higher number of time series, making their results more stable and less prone to overfitting compared to smaller datasets. We evaluate all methods across multiple prediction horizons $H\in{96,192,336,720}$ for standard benchmarks and $H\in{24,36,48,60}$ for the ILI dataset.
4.2.2 Imputation
The performance of imputation is assessed on benchmark datasets from the ETT [zhou2021informer], Electricity [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather] domains, which commonly encounter missing data in real-world applications. Following the protocol established by TimesNet [wu2023timesnettemporal2dvariationmodeling], we evaluated performance under random masking ratios of $12.5\%,25\%,37.5\%,50\%$}. The evaluation includes recent SOTA baselines such as GPT2(3) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
4.2.3 Classification
For classification tasks, we utilized ten multivariate benchmark datasets from the UEA time series classification archive [bagnall2018ueamultivariatetimeseries]. These datasets encompass diverse applications including gesture and action recognition, heartbeat-based medical diagnosis, and other real-world scenarios. The evaluation compares with recent SOTA baselines such as GPT2(6) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], DLinear [zeng2022transformerseffectivetimeseries], ETSFormer [woo2022etsformer], FEDformer [zhou2022fedformer], and Informer [zhou2021informer].
4.2.4 SOTA Baselines for Generalizability Study
To rigorously evaluate the generalizability of the proposed MSTN framework, we conducted cross-domain benchmarking across seven publicly available international datasets. The evaluation encompasses datasets from diverse domains: human safety, with the Boubézoul et al. [BOUBEZOUL2020105577] (France) fall event dataset and the Rodegast et al. [DARUS33012023] (Germany) simulator for risky riding behaviour and collisions; general human activity, using UCI-HAR [har240] (Italy) and PAMAP2 [pamap2_physical_activity_monitoring_231] (USA) to assess robustness in physical activity recognition; animal behavior monitoring via the ActBe-Calf dataset [DISSANAYAKE2025111462] (Ireland) for welfare assessment; and mechanical systems, including MetroPT-3 [metropt3] (Portugal) for predictive maintenance and the NASA Turbofan Engine [NasaTurboFan] (USA) for sensor degradation forecasting. This selection tests model transferability across healthcare, activity recognition, agricultural technology, and industrial monitoring. All models were evaluated under a consistent five-fold cross-validation framework to ensure a robust and fair comparison, including recent architectures such as TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
4.3 Training and Model Configuration
The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1+cu118 and trained on an NVIDIA T400 GPU (4 GB VRAM). The architecture processes variable-length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128→64 filters). Training uses AdamW [kingma2017adammethodstochasticoptimization] optimizer with learning rate $3\times 10^{-4}$, batch size 64, and task-specific objectives (Focal Loss for classification, MSE for forecasting and imputation). Models train for up to 100 epochs with early stopping based on validation performance for task. Dataset-specific configurations include 96-step lookback for Traffic data and 50-step sequences for physical therapy datasets, with min-max feature normalization applied throughout.
5 Acknowledgments
This research is supported by the IIT Indore Young Faculty Research Catalysing Grant (YFRCG) Scheme ( Project ID: IITI/YFRCG/2023-24/01), Indian Institute of Technology Indore, India.
6 Data Availability
All datasets used in this study are publicly available from their respective sources as cited throughout the paper.
7 Author Contributions
Sumit S Shevtekar led the research design, developed the MSTN architecture, implemented all software, conducted experiments across classification, forecasting and imputation tasks, performed data analysis and visualization, and wrote and edited the original draft. Chandresh K Maurya contributed to methodology design, experimental validation, and manuscript writing, review and editing. Gourab Sil participated in manuscript review and editing. All authors reviewed and approved the final manuscript.
8 Competing Interests
The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.

### v2

@@
-MSTN: Fast and Efficient Multivariate Time Series Model
+MSTN: Fast and Efficient Multivariate Time Series Prediction Model
 Abstract
-Real-world time-series data is highly non stationary and complex in dynamics that operate across multiple timescales, ranging from fast, short-term changes to slow, long-term trends.
-Most existing models rely on fixed-scale structural priors, such as patch-based tokenization, fixed frequency transformations, or frozen backbone architectures.
-This often leads to over-regularization of temporal dynamics, which limits their ability to adaptively model the full spectrum of temporal variations and impairs their performance on unpredictable, Sudden, high-magnitude events.
-To address this, we introduce the Multi-scale Temporal Network (MSTN), a novel deep learning architecture founded on a hierarchical multi-scale and sequence modeling principle.
-The MSTN framework integrates three core components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid for local patterns (ii) a sequence modeling component for long-range temporal dependencies.
-We empirically validate this with Bidirectional Long Short-Term Memory (LSTM) and Transformer variants, establishing a flexible foundation for future architectural advancements.
-and (iii) a gated fusion mechanism augmented with squeeze-and-excitation (SE) and multi-head temporal attention (MHTA) for dynamic, context-aware feature integration.
-This design enables MSTN to adaptively model temporal patterns from milliseconds to long-range dependencies within a unified framework.
-Extensive evaluations across time-series long-horizon forecasting, imputation, classification and generalizability study demonstrate that MSTN achieves competitive state-of-the-art (SOTA) performance, showing improvements over contemporary approaches including EMTSF, LLM4TS, HiMTM, TIME-LLM, MTST, SOFTS, iTransformer, TimesNet, and PatchTST.
-In total, MSTN establishes new SOTA performance on 24 of 32 benchmark datasets, demonstrating its consistent performance across diverse temporal tasks.
+Real-world time series often exhibit strong non-stationarity, complex nonlinear dynamics, and behaviour expressed across multiple temporal scales, from rapid local fluctuations to slow-evolving long-range trends.
+However, many contemporary architectures impose rigid, fixed-scale structural priors—such as patch-based tokenization, predefined receptive fields, or frozen backbone encoders—which can over-regularize temporal dynamics and limit adaptability to abrupt high-magnitude events.
+To handle this, we introduce the Multi-scale Temporal Network (MSTN), a hybrid neural architecture grounded in an Early Temporal Aggregation principle.
+MSTN integrates three complementary components: (i) a multi-scale convolutional encoder that captures fine-grained local structure; (ii) a sequence modeling module that learns long-range dependencies through either recurrent or attention-based mechanisms; and (iii) a self-gated fusion stage incorporating squeeze–excitation and multi-head attention to dynamically modulate cross-scale representations.
+This design enables MSTN to flexibly model temporal patterns spanning milliseconds to extended horizons, while avoiding the computational burden typically associated with long-context models.
+Across extensive benchmarks covering forecasting, imputation, classification, and cross-dataset generalization, MSTN consistently delivers state-of-the-art performance, outperforming recent leading approaches including TIME-LLM, HiMTM, SOFTS, LLM4TS, TimesNet, and PatchTST, and establishing new best results on 24 out of 32 datasets.
+Despite its strong performance, MSTN remains lightweight and supports fast inference, making it well suited for deployment on edge devices and resource-constrained environments.
 keywords:
@@
 [inst1]organization=Department of Computer Science and Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
-[…]
-[inst2]organization=Department of Civil Engineering, Indian Institute of Technology Indore, addressline=Indore, postcode=452020, state=Madhya Pradesh, country=India
 1 Introduction
-Multivariate time series analysis is a crucial component of modern AI, with significant real-world applications ranging from healthcare monitoring and industrial predictive maintenance to behavioral analytics.
+Multivariate time series (MTS) analysis has been a classical problem with significant real-world applications including healthcare monitoring [10.1145/3531326], traffic management [kadiyala2014multivariate], environment [gruca2022weather4cast], industrial predictive maintenance, energy  [6714975] and behavioral analytics.
 It has been extensively studied for many years and has been widely used in tasks such as weather forecasting, imputing missing data for data mining, and classification of trajectories for action recognition [wu2022autoformer, zhao2024himtmhierarchicalmultiscalemasked].
@@
 This fundamental characteristic forces models to extract information from temporal variation patterns of continuity, periodicity, and trend, making robust temporal modeling exceptionally challenging [wu2023timesnettemporal2dvariationmodeling, lim2021time].
-Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving SOTA results [nie2023timeseriesworth64].
+Deep learning (DL) models have demonstrated strong performance not only in forecasting tasks, but also in representation learning, where abstract features can be extracted and transferred to downstream tasks such as classification and forecasting, achieving state-of-the-art (SOTA) results [nie2023timeseriesworth64].
 Recent findings show that simple linear models can outperform sophisticated Transformer variants on common benchmarks [zeng2022transformerseffectivetimeseries], highlighting fundamental limitations in contemporary architectures.
-This is particularly critical in safety-sensitive domains like behavioral analytics and industrial monitoring, where patterns are often non-periodic and evolve across multiple timescales.
 The design of temporal models faces the constant challenge of balancing model strength with computational efficiency.
-Models like recurrent neural networks (RNN) and their gated variants [6795963, lai2018modelinglongshorttermtemporal] offer a strong sequential modeling but suffer from vanishing gradients and limited parallelism.
+Models such as convolutional neural networks (CNNs) [franceschi2019unsupervised, bai2018empirical] and recurrent neural networks (RNNs) with their gated variants [6795963, lai2018modelinglongshorttermtemporal] suffer from vanishing gradients and limited parallelism, which has shifted the paradigm towards more complex models.
 Temporal convolutional networks (TCN)  [he2019temporal] improve computational efficiency but are constrained by fixed receptive fields.
@@
 These challenges have shaped several predominant, yet ultimately constrained, modern paradigms.
-One approach, exemplified by PatchTST [nie2023timeseriesworth64], adapts the patching technique from Vision Transformers, grouping adjacent time points into sub-series tokens.
-This approach successfully captures local semantics and benefits from channel-independent processing, but its reliance on a fixed patch size creates a fundamental rigidity, locking the model into a single, pre-defined scale of analysis.
-Another paradigm, improved by TimesNet [wu2023timesnettemporal2dvariationmodeling], reframes the problem by transforming 1D time series into 2D tensors based on identified periods.
-While powerful for data with strong periodic components, this approach is intrinsically dependent on reliable periodicity detection and the 2D transformation prior, making it less suited for non-stationary, irregular, or non-periodic patterns where such clear structure is absent.
-Building on these concepts, MTST [zhang2024multiresolutiontimeseriestransformerlongterm] employs a multi-resolution patch-based design with a multi-branch architecture for simultaneous modeling of diverse temporal patterns at different resolutions.
-However, its fixed choice of resolutions is less flexible and may fail to generalize across datasets with irregular or heterogeneous temporal patterns.
-Patch tokenization can also lose fine-grained dynamics, while the reliance on relative positional encoding restricts robustness in non-seasonal or event-driven sequences.
-We identify a fundamental limitation on architectures that impose specific single-scale structural constraints, such as a fixed patch size, a period-based 2D shape, or frozen backbone architectures.
-These prior models struggle with the inherently multi-scale nature of many real-world temporal incidents.
-Their architecture limits their ability to model discriminative patterns that manifest simultaneously across different timescales, from brief micro-events to slow macro and long-term trends, such as found in risky driving behaviour.
-This gap highlights the need for a more flexible temporal representation that can primarily adapt to this hierarchical structure without relying on a single, potentially misaligned, structural assumption.
-To address these challenges, we propose a DL model (MSTN) that naturally captures temporal hierarchies without relying on fixed and repeated structural assumptions.
-MSTN unifies multi-scale feature learning with global sequence modeling through a dual-path architecture with a sequence modeling core.
-We empirically validate this design with two instantiations: i) MSTN–BiLSTM, which excels at modeling local continuity and behavioral smoothness, and ii) MSTN–Transformer, which captures local complex patterns and long-range contextual dependencies.
-In both variants, convolutional encoders extract fine-grained local features, while the temporal core captures long-range dependencies.
-Their outputs are integrated via a gated fusion mechanism enhanced with SE recalibration and MHTA, yielding compact, discriminative, and temporally coherent representations.
-This design provides an unprecedented trade-off between representational capacity and computational efficiency.
+These can be formalized into three primary Channel Strategies (CS) [qiu2025comprehensivesurveydeeplearning]: Channel Dependence (CD), which favors high feature completeness at a computational cost; Channel Independence (CI), which prioritizes speed by isolating channels; and Channel Partiality (CP), which aims for a balance by reducing channel complexity.
+CI based models like PatchTST [nie2023timeseriesworth64] improves local semantic modeling through patch tokenization but introduces rigidity through fixed patch sizes; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures.
+CD based models such as TimesNet [wu2023timesnettemporal2dvariationmodeling] transforms time series into periodic 2D structures but depends heavily on accurate period discovery, limiting robustness for irregular or non-stationary signals.
+MTST [zhang2024multiresolutiontimeseriestransformerlongterm] models multiple resolutions but relies on fixed resolution sets and may dismiss fine-grained dynamics.
+Even highly efficient recent systems such as SOFTS [han2024softsefficientmultivariatetime] prioritize channel efficiency through linear factorization, inherently sacrificing temporal completeness and weakening multi-scale temporal representation.
+Similarly, the CP Channels strategy in MCformer [10533212] achieves efficiency by partially blending a specific number of channels, but this partiality limits the recovery of full inter-channel context.
+Consequently, such architectures struggle to sustain fidelity in complex long-sequence environments.
+We identify a fundamental limitation on these approaches: the imposition of single-scale architectural assumptions such as fixed patch sizes, fixed periodicity-driven 2D mappings, struggle to capture discriminative cues manifesting across diverse temporal horizons, or static backbone structures.
+However, real-world temporal phenomena are inherently multi-scale: micro-events, meso-scale fluctuations, and macro trends often coexist.
+Consequently, existing models struggle to capture discriminative cues manifesting across diverse temporal horizons, particularly in domains such as risky driving behaviour, physiological safety monitoring, or mechanical prognostics—where multi-scale variability is intrinsic.
+This gap motivates a more flexible, hierarchy-aware temporal representation without relying on a single predefined structure.
+To address these limitations, we propose the Multi-scale Temporal Network (MSTN), a hybrid architecture founded on the Early Temporal Aggregation (ETA) principle.
+MSTN naturally captures hierarchical temporal patterns without fixed structural assumptions through its central innovation: ETA condenses the full temporal input into a single static feature vector via an $L\rightarrow 1$ transformation, effectively neutralizing the $\mathcal{O}(L^{2})$ inference bottleneck while preserving global completeness.
+This design allows the dual encoder to retain $\mathcal{O}(L^{2})$ modeling capacity during representation extraction, while all subsequent refinement operates in guaranteed constant-time $\mathcal{O}(1)$ complexity.
+MSTN unifies multi-scale feature extraction with global temporal reasoning through a dual-path architecture: convolutional encoders capture fine-grained local structure, while a Bidirectional Long Short-Term Memory (BiLSTM) or Transformer backbone models long-range dependencies.
+Their outputs are fused through a self-gated fusion (SGF) mechanism enriched by squeeze-and-excitatio (SE) recalibration and multi-head attention (MHA), producing compact, discriminative, and temporally coherent representations.
+We empirically validate the MSTN framework through two realizations: MSTN–BiLSTM, ideal for smooth continuous behaviors, and MSTN–Transformer, suitable for complex, long-range event modeling.
 We demonstrate that MSTN establishes new SOTA performance on 24 of 32 benchmark datasets for time series analysis.
-This includes 9 forecasting benchmarks, and 6 imputation tasks and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications.
+This includes 9 forecasting benchmarks, 6 imputation tasks, and 17 classification datasets (10 from the standard UEA archive and 7 additional international benchmarks), demonstrating its versatility and improved performance across diverse applications.
 MSTN establishes a new SOTA in long-term forecasting, attaining the lowest prediction error on 6 out of 9 standard benchmarks and consistently exceeds leading architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], GPT2(6) [zhou2023fitsallpowergeneraltime], SOFTS [han2024softsefficientmultivariatetime] iTransformer [liu2024itransformerinvertedtransformerseffective], and PatchTST [nie2023timeseriesworth64].
 Complementing these results, the MSTN-Transformer variant achieves first-place rankings 31 out of 48 ($\approx$65%) in all 48 imputation scenarios across the ETT, Electricity, and Weather datasets, demonstrating strong resilience under up to 50% missingness.
-evaluation scenarios and secure the highest average score on 5 out of 6 datasets.
 In the classification task, MSTN achieves improved performance over recent SOTA architectures such as GTP2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], LigthtTS  [zhang2022morefastmultivariatetime] and FEDformer [zhou2022fedformer].
 We further evaluate MSTN for generalizability across 7 benchmark datasets from various domains like human safety, physical activity recognition, animal-welfare monitoring, and mechanical prognostics.
-Without domain-specific tuning, surprisingly, MSTN achieves uniformly strong performance, ranks first across all datasets, establishing new benchmarks and illustrating its robustness under diverse temporal dynamics.
-Finally, MSTN-Transformer achieves unprecedented performance with 0.020 MSE, 2.87 MB footprint and 0.72 ms inference time, representing a 22.4× accuracy improvement and 34.7× speedup over SOFTS [han2024softsefficientmultivariatetime].
-The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.155 ms) significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al.
+Without domain-specific tuning, MSTN achieves improved performance, ranking first across all datasets, and illustrates its robustness under diverse temporal dynamics.
+Concretely, MSTN-Transformer achieves unprecedented performance with 2.87 MB memory footprint, and 0.72 ms inference latency, representing a $22.4\times$ accuracy improvement and $34.7\times$ speedup over SOFTS [han2024softsefficientmultivariatetime].
+The model further demonstrates improved classification accuracy (99.53%) with ultra-low latency (0.15 ms), significantly outperforming TimesNet (99.18%, 5.07 ms) and PatchTST (98.64%, 4.80 ms) on the Rodegast et al.
 [DARUS33012023] dataset.
-While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN-Transformer as an ideal solution for real-time and edge-AI temporal modeling applications.
-Our comprehensive evaluation establishes MSTN as a new benchmark, achieving SOTA performance on 24 of 32 datasets across forecasting, imputation, and classification tasks.
+While maintaining compact model sizes (0.54–7.14 MB), this balance of SOTA accuracy, robustness, and efficiency positions MSTN as a preferred solution for real-time and edge-AI temporal modeling applications.
 2 Results
-This section presents a comprehensive evaluation of the proposed MSTN.
-Demonstrating its SOTA performance in the long-term forecasting, imputation, classification, and generalizability study.
-MSTN (MSTN-Transformer and MSTN-BiLSTM) achieves competitive performance, showing improvements over recent DL, Transformer based, and time series approaches.
+This section presents a comprehensive evaluation of the proposed MSTN model.
+We demonstrate its performance on three time-series tasks: (1) long-term forecasting, (2) imputation, (3) classification.
+Additionally, we also present the generalizability study on standard time-series benchmark datasets.
+2.1 Benchmark Datasets
+Table 1: Benchmark datasets and imputation performance comparison.
+(a) Benchmark datasets for forecasting, imputation, and classification (10 UEA).
+Task Datasets Features Timesteps Metrics Length
+Imputation ETTh1/ETTh2 7 17,420 MSE, MAE 96
+ETTm1/ETTm2 7 69,680
+Electricity 321 26,304
+Weather 21 52,696
+Forecasting ETTh1/ETTh2 7 17,420 MSE, MAE 96–720
+ETTm1/ETTm2 7 69,680
+Electricity 321 26,304
+Traffic 862 17,544
+Weather 21 52,696
+ILI 7 966
+Classification EthanolConcentration 3 1,751 Accuracy 29-1,751
+FaceDetection 144 62
+Handwriting 3 152
+Heartbeat 61 405
+JapaneseVowels 12 29
+PEMS-SF 963 144
+SelfRegulationSCP1 6 896
+SelfRegulationSCP2 7 1,152
+SpokenArabicDigits 13 93
+UWaveGestureLibrary 3 315
+(b) Imputation task results. Random masking ratios of 12.5%, 25%, 37.5%, and 50% are applied. Red and blue denote first and second best results, respectively.
+Models MSTN-Tra. MSTN BiL. GPT2(3) TimesNet PatchTST LightTS DLinear Stationary
+(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
+Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
+ETTh1
+12.5% 0.024 0.040 0.032 0.047 0.043 0.140 0.057 0.159 0.093 0.201 0.240 0.345 0.151 0.267 0.060 0.165
+25% 0.050 0.083 0.067 0.096 0.054 0.156 0.069 0.178 0.107 0.217 0.265 0.364 0.180 0.292 0.080 0.189
+37.5% 0.083 0.131 0.103 0.147 0.072 0.180 0.084 0.196 0.120 0.230 0.296 0.382 0.215 0.318 0.102 0.212
+50% 0.127 0.187 0.148 0.204 0.107 0.216 0.102 0.215 0.141 0.248 0.334 0.404 0.257 0.347 0.133 0.240
+Avg 0.071 0.110 0.088 0.124 0.069 0.173 0.078 0.187 0.115 0.224 0.284 0.373 0.201 0.306 0.094 0.201
+ETTh2
+12.5% 0.009 0.024 0.022 0.040 0.039 0.125 0.040 0.130 0.057 0.152 0.101 0.231 0.100 0.216 0.042 0.133
+25% 0.022 0.054 0.043 0.078 0.044 0.135 0.046 0.141 0.061 0.158 0.115 0.246 0.127 0.247 0.049 0.147
+37.5% 0.038 0.089 0.069 0.121 0.051 0.147 0.052 0.151 0.067 0.166 0.126 0.257 0.158 0.276 0.056 0.158
+50% 0.066 0.135 0.102 0.170 0.059 0.158 0.060 0.162 0.073 0.174 0.136 0.268 0.183 0.299 0.065 0.170
+Avg 0.034 0.076 0.059 0.102 0.048 0.141 0.049 0.146 0.065 0.163 0.119 0.250 0.142 0.259 0.053 0.152
+ETTm1
+12.5% 0.024 0.038 0.036 0.049 0.017 0.085 0.019 0.092 0.041 0.130 0.075 0.180 0.058 0.162 0.026 0.107
+25% 0.050 0.080 0.072 0.099 0.022 0.096 0.023 0.101 0.044 0.135 0.093 0.206 0.080 0.193 0.032 0.119
+37.5% 0.082 0.124 0.112 0.150 0.029 0.111 0.029 0.111 0.049 0.143 0.113 0.231 0.103 0.219 0.039 0.131
+50% 0.121 0.176 0.160 0.210 0.040 0.128 0.036 0.124 0.055 0.151 0.134 0.255 0.132 0.248 0.047 0.145
+Avg 0.069 0.104 0.095 0.127 0.028 0.105 0.027 0.107 0.047 0.140 0.104 0.218 0.093 0.206 0.036 0.126
+ETTm2
+12.5% 0.039 0.048 0.052 0.057 0.017 0.076 0.018 0.080 0.026 0.094 0.034 0.127 0.062 0.166 0.021 0.088
+25% 0.080 0.097 0.107 0.116 0.020 0.080 0.020 0.085 0.028 0.099 0.042 0.143 0.085 0.196 0.024 0.096
+37.5% 0.132 0.153 0.164 0.176 0.022 0.087 0.023 0.091 0.030 0.104 0.051 0.159 0.106 0.222 0.027 0.103
+50% 0.187 0.215 0.225 0.241 0.025 0.095 0.026 0.098 0.034 0.110 0.059 0.174 0.131 0.247 0.030 0.108
+Avg 0.109 0.128 0.137 0.147 0.021 0.084 0.022 0.088 0.029 0.102 0.046 0.151 0.096 0.208 0.026 0.099
+ECL
+12.5% 0.002 0.014 0.034 0.048 0.080 0.194 0.085 0.202 0.055 0.160 0.102 0.229 0.092 0.214 0.093 0.210
+25% 0.005 0.028 0.069 0.096 0.087 0.203 0.089 0.206 0.065 0.175 0.121 0.252 0.118 0.247 0.097 0.214
+37.5% 0.008 0.044 0.108 0.147 0.094 0.211 0.094 0.213 0.076 0.189 0.141 0.273 0.144 0.276 0.102 0.220
+50% 0.013 0.063 0.150 0.202 0.101 0.220 0.100 0.221 0.091 0.208 0.160 0.293 0.175 0.305 0.108 0.228
+Avg 0.007 0.037 0.090 0.123 0.090 0.207 0.092 0.210 0.072 0.183 0.131 0.262 0.132 0.260 0.100 0.218
+Weather
+12.5% 0.003 0.016 0.026 0.036 0.026 0.049 0.025 0.045 0.029 0.049 0.047 0.101 0.039 0.084 0.027 0.051
+25% 0.007 0.032 0.051 0.070 0.028 0.052 0.029 0.052 0.031 0.053 0.052 0.111 0.048 0.103 0.029 0.056
+37.5% 0.011 0.049 0.078 0.109 0.033 0.060 0.031 0.057 0.035 0.058 0.058 0.121 0.057 0.117 0.033 0.062
+50% 0.015 0.068 0.111 0.153 0.037 0.065 0.034 0.062 0.038 0.063 0.065 0.133 0.066 0.134 0.037 0.068
+Avg 0.009 0.041 0.066 0.092 0.031 0.056 0.030 0.054 0.033 0.056 0.055 0.117 0.052 0.110 0.032 0.059
+We evaluate the proposed method across three fundamental time series tasks: forecasting, missing value imputation, and classification using established public benchmark datasets.
+Table 1 provides a detailed overview of all datasets.
+Missing value imputation is performed under a Missing Completely at Random (MCAR) setting, with masking ratios ranging from 12.5%, 25%, 37.5% and 50%.
+Benchmarks include ETTh1, ETTh2, ETTm1, ETTm2, Electricity, and Weather datasets [zhou2021informer, electricityloaddiagrams20112014_321, Kolle2025_BGCJenaWeather].
+Imputation results are quantified using Mean Squared Error (MSE) and Mean Absolute Error (MAE) over sequence lengths of 96 time steps.
+Long-term forecasting performance is assessed on nine widely-used benchmarks: ETTh1, ETTh2 (Electricity Transformer Temperature datasets of the China State Grid, recorded at an hourly resolution), ETTm1, ETTm2 (15-minute variants of the same transformer monitoring system), Electricity (ECL - hourly household consumption data from 321 U.S.
+clients), Traffic (hourly road occupancy rates from CalTrans loop detectors in the United States), Weather (10-minute meteorological observations from 21 U.S.
+stations), Exchange (daily currency exchange rates across major global economies), and ILI (weekly influenza-like illness records from the U.S.
+CDC).
+These datasets are extensively used for time-series benchmarking and are publicly available [wu2022autoformer].
+Forecasting horizons are set between 96, 192, 336, and 720 time steps.
+Long-term forecasting performance is evaluated using MSE and MAE.
+The multivariate time series classification is evaluated on ten diverse datasets from the University of East Anglia (UEA) archive [bagnall2018ueamultivariatetimeseries], which includes applications from healthcare to activity recognition.
+To rigorously assess cross-domain transferability, we utilize seven publicly available international datasets spanning diverse applications: Human Safety/Healthcare (Fall Event [BOUBEZOUL2020105577], Risky Driving [DARUS33012023]); Human Activity Recognition (HAR) (UCI-HAR [har240], PAMAP2 [pamap2_physical_activity_monitoring_231]); Agricultural Technology/Welfare (ActBe-Calf [DISSANAYAKE2025111462]); and Industrial Monitoring/Predictive Maintenance (MetroPT-3 [metropt3], NASA Turbofan Engine [NasaTurboFan]).
+2.2 Imputation
+Imputation is the task of filling in missing values in a dataset.
+In real-world time series data, sensors can fail or data can be corrupted, which may create gaps in the data.
+This task tests the ability of the model to intelligently reconstruct missing data based on the surrounding context, which is vital for maintaining data quality.
+As presented in Table 1, MSTN-Transformer achieves SOTA performance in 32 out of 48 ($\approx$66%) imputation evaluation scenarios and delivers the improved average results on 5 of the 6 benchmark datasets, while MSTN-BiLSTM demonstrates competitive performance.
+Following the TimesNet benchmark protocol [wu2023timesnettemporal2dvariationmodeling], we evaluate imputation performance on electricity and weather domain datasets, including ETT [zhou2021informer], ECL [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather].
+To simulate real-world scenarios, we employ random masking ratios of {12.5%, 25%, 37.5%, 50%}.
+This shows improved performance over several recent approaches, including GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
 2.3 Long-Term Forecasting
-Table 1: Multivariate forecasting results with prediction lengths $H\in\{96,192,336,720\}$ for others and for ILI $H\in\{24,36,48,60\}$. Red/Blue: First/Second ranks.
+Table 2: Multivariate long-term forecasting results.
+Input lookback window $L=96$ ($L=36$ for ILI).
+Prediction horizons $H\in\{96,192,336,720\}$ ($H\in\{24,36,48,60\}$ for ILI).
+Avg is averaged from all four prediction lengths.
+Red/Blue: First/Second ranks.
 Models MSTN-Trans. MSTN BiL. EMTSF LLM4TS HiMTM TIME-LLM MTST SOFTS iTransf.
@@
 Avg 0.165 0.302 0.166 0.332 1.588 0.745 - - - - 1.436 0.801 - - - - - -
-We evaluate long-term forecasting performance using established protocols from the prior literature [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64] across nine diverse benchmark datasets.
+Time series forecasting plays a crucial role in applications such as weather prediction, traffic management, and energy consumption planning.
+Long-term forecasting aims to predict future values far ahead in time.
+For example, it involves using historical weather data to estimate not only the temperature of tomorrow but also the conditions over the coming days or even weeks.
+We evaluate long-term forecasting performance using established protocols from the prior literature  [nie2023timeseriesworth64, wu2023timesnettemporal2dvariationmodeling, wu2022autoformer] across nine diverse benchmark datasets.
 The prediction horizons are set to $H\in\{96,192,336,720\}$ for the standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI data set.
-As presented in Table 1, the MSTN-Transformer and MSTN BiLSTM shows improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm],SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
+As presented in Table 2, MSTN-Transformer and MSTN-BiLSTM show improved performance compared to specialized time series forecasting architectures, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
 […]
-2.2 Imputation
-Table 2: Imputation Task Results Comparison. We randomly mask 12.5%, 25%, 37.5%, and 50% time points to compare the model performance under different missing degrees. Red/Blue: First/Second ranks.
-Models MSTN-Tra. MSTN BiL. GPT2(3) TimesNet PatchTST LightTS DLinear Stationary
-(Ours) (Ours) (2023) (2023) (2023) (2022) (2023) (2022)
-Mask R. MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
-ETTh1
-12.5% 0.024 0.040 0.032 0.047 0.043 0.140 0.057 0.159 0.093 0.201 0.240 0.345 0.151 0.267 0.060 0.165
-25% 0.050 0.083 0.067 0.096 0.054 0.156 0.069 0.178 0.107 0.217 0.265 0.364 0.180 0.292 0.080 0.189
-37.5% 0.083 0.131 0.103 0.147 0.072 0.180 0.084 0.196 0.120 0.230 0.296 0.382 0.215 0.318 0.102 0.212
-50% 0.127 0.187 0.148 0.204 0.107 0.216 0.102 0.215 0.141 0.248 0.334 0.404 0.257 0.347 0.133 0.240
-Avg 0.071 0.110 0.088 0.124 0.069 0.173 0.078 0.187 0.115 0.224 0.284 0.373 0.201 0.306 0.094 0.201
-ETTh2
-12.5% 0.009 0.024 0.022 0.040 0.039 0.125 0.040 0.130 0.057 0.152 0.101 0.231 0.100 0.216 0.042 0.133
-25% 0.022 0.054 0.043 0.078 0.044 0.135 0.046 0.141 0.061 0.158 0.115 0.246 0.127 0.247 0.049 0.147
-37.5% 0.038 0.089 0.069 0.121 0.051 0.147 0.052 0.151 0.067 0.166 0.126 0.257 0.158 0.276 0.056 0.158
-50% 0.066 0.135 0.102 0.170 0.059 0.158 0.060 0.162 0.073 0.174 0.136 0.268 0.183 0.299 0.065 0.170
-Avg 0.034 0.076 0.059 0.102 0.048 0.141 0.049 0.146 0.065 0.163 0.119 0.250 0.142 0.259 0.053 0.152
-ETTm1
-12.5% 0.024 0.038 0.036 0.049 0.017 0.085 0.019 0.092 0.041 0.130 0.075 0.180 0.058 0.162 0.026 0.107
-25% 0.050 0.080 0.072 0.099 0.022 0.096 0.023 0.101 0.044 0.135 0.093 0.206 0.080 0.193 0.032 0.119
-37.5% 0.082 0.124 0.112 0.150 0.029 0.111 0.029 0.111 0.049 0.143 0.113 0.231 0.103 0.219 0.039 0.131
-50% 0.121 0.176 0.160 0.210 0.040 0.128 0.036 0.124 0.055 0.151 0.134 0.255 0.132 0.248 0.047 0.145
-Avg 0.069 0.104 0.095 0.127 0.028 0.105 0.027 0.107 0.047 0.140 0.104 0.218 0.093 0.206 0.036 0.126
-ETTm2
-12.5% 0.039 0.048 0.052 0.057 0.017 0.076 0.018 0.080 0.026 0.094 0.034 0.127 0.062 0.166 0.021 0.088
-25% 0.080 0.097 0.107 0.116 0.020 0.080 0.020 0.085 0.028 0.099 0.042 0.143 0.085 0.196 0.024 0.096
-37.5% 0.132 0.153 0.164 0.176 0.022 0.087 0.023 0.091 0.030 0.104 0.051 0.159 0.106 0.222 0.027 0.103
-50% 0.187 0.215 0.225 0.241 0.025 0.095 0.026 0.098 0.034 0.110 0.059 0.174 0.131 0.247 0.030 0.108
-Avg 0.109 0.128 0.137 0.147 0.021 0.084 0.022 0.088 0.029 0.102 0.046 0.151 0.096 0.208 0.026 0.099
-ECL
-12.5% 0.002 0.014 0.034 0.048 0.080 0.194 0.085 0.202 0.055 0.160 0.102 0.229 0.092 0.214 0.093 0.210
-25% 0.005 0.028 0.069 0.096 0.087 0.203 0.089 0.206 0.065 0.175 0.121 0.252 0.118 0.247 0.097 0.214
-37.5% 0.008 0.044 0.108 0.147 0.094 0.211 0.094 0.213 0.076 0.189 0.141 0.273 0.144 0.276 0.102 0.220
-50% 0.013 0.063 0.150 0.202 0.101 0.220 0.100 0.221 0.091 0.208 0.160 0.293 0.175 0.305 0.108 0.228
-Avg 0.007 0.037 0.090 0.123 0.090 0.207 0.092 0.210 0.072 0.183 0.131 0.262 0.132 0.260 0.100 0.218
-Weather
-12.5% 0.003 0.016 0.026 0.036 0.026 0.049 0.025 0.045 0.029 0.049 0.047 0.101 0.039 0.084 0.027 0.051
-25% 0.007 0.032 0.051 0.070 0.028 0.052 0.029 0.052 0.031 0.053 0.052 0.111 0.048 0.103 0.029 0.056
-37.5% 0.011 0.049 0.078 0.109 0.033 0.060 0.031 0.057 0.035 0.058 0.058 0.121 0.057 0.117 0.033 0.062
-50% 0.015 0.068 0.111 0.153 0.037 0.065 0.034 0.062 0.038 0.063 0.065 0.133 0.066 0.134 0.037 0.068
-Avg 0.009 0.041 0.066 0.092 0.031 0.056 0.030 0.054 0.033 0.056 0.055 0.117 0.052 0.110 0.032 0.059
-As presented in Table 2, MSTN-Transformer achieves SOTA performance in 32 out of 48 ($\approx$66%) imputation evaluation scenarios and delivers the improved average results on 5 of the 6 benchmark datasets, while MSTN-BiLSTM demonstrates competitive performance.
-Following the TimesNet benchmark protocol [wu2023timesnettemporal2dvariationmodeling], we evaluate imputation performance on electricity and weather domain datasets, including ETT [zhou2021informer], ECL [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather].
-To simulate real-world scenarios, we employ random masking ratios of 12.5%, 25%, 37.5%, 50%.
-This shows improved performance over several recent approaches, including GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
 2.4 Time-Series Classification
-Table 3 presents the classification results across 10 UEA standard benchmark datasets.
-The proposed MSTN framework is compared against SOTA time-series classification methods, including GPT-2(6) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime].
-MSTN demonstrates competitive performance in multiple datasets.
-MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
 Table 3: Comprehensive evaluation of MSTN across diverse time-series benchmarks.
@@
 Dataset Det. Dom. Cnt. A.Block Acc.(%) F1 Prec. Rec. Size I.Time Acc.(%) I.Time Acc.(%) I.Time Prior Acc.
-Rodeg.[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
+Rodegast[DARUS33012023] Sim. Hum. DE. BiLSTM 99.20 0.992 0.993 0.991 4.39 0.81 99.18 5.07 98.64 4.80 91.00[RF,GB][Rodegast2024]
 Transf. 99.53 0.947 0.951 0.942 0.54 0.155
-Boubez.[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
+Boubezoul[BOUBEZOUL2020105577] Real. Hum. FR. BiLSTM 93.75 0.945 0.940 0.950 4.16 2.79 93.00 4.20 91.37 4.68 91.59[DT][fi15100333]
 Transf. 91.76 0.927 0.921 0.934 0.96 0.24
@@
 Transformer), I.Time: Inference time, †RMSE values.
+Classification involves categorizing entire time series sequences.
+It is used to identify which type of event or pattern the sequence represents.
+For example, this could be the classification of a segment of sensor data as walking, running, falling, or collision activity, which is fundamental for automated monitoring and diagnosis systems.
+Performance is evaluated using the standard metric of classification accuracy.
+Table 3 presents the classification results across 10 UEA standard benchmark datasets.
+The proposed MSTN framework is compared against SOTA time-series classification methods, including TimesNet [wu2023timesnettemporal2dvariationmodeling], GPT-2(6) [zhou2023fitsallpowergeneraltime], FEDformer [zhou2022fedformer], and LightTS [zhang2022morefastmultivariatetime].
+MSTN demonstrates competitive performance in multiple datasets.
+MSTN-Transformer achieves perfect classification (100%) on JapaneseVowels and near-perfect performance on SpokenArabic (99.32%), while MSTN-BiLSTM achieves improved accuracy on Heartbeat (81.49%), JapaneseVowels (100%) and PEMS-SF (96.59%).
 2.5 Generalizability Study
-As summarized in Table 3, the proposed MSTN framework demonstrates remarkable generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics.
-Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64] in both accuracy and computational efficiency.
+A robust model should not be a specialist in just one area; it should adapt to new tasks and data distributions across different data and domains.
+This experiment evaluates the ability of the MSTN to transfer its knowledge across diverse domains, from human activity recognition to mechanical fault prediction.
+As summarized in Table 3, the proposed MSTN framework demonstrates generalization in seven international benchmark data sets such as human safety, activity recognition, animal welfare, and mechanical prognostics.
+Without domain-specific tuning, MSTN shows improved performance over the latest time series SOTA models, including TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
 MSTN achieves competitive performance across fundamentally different domains, demonstrating its capabilities as a universal temporal model.
@@
 The framework also excels in real-world fall detection with 93.75% accuracy, surpassing dedicated DT-based methods.
-MSTN-BiLSTM also shows strong performance in animal welfare[DISSANAYAKE2025111462] assessment with 93.00% accuracy on calf behavior recognition.
-Furthermore, MSTN-Transformer demonstrates exceptional capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan dataset[NasaTurboFan] and significantly outperforming all baseline predictors in remaining useful life estimation.
-MSTN delivers this performance with good computational efficiency.
-For instance, on the Rodegast [DARUS33012023] Germany data set, the MSTN-Transformer variant achieves 0.155 ms/sample inference latency 33× faster than TimesNet and 31.0× faster than PatchTST while maintaining compact model size.
-This efficiency enables deployment on resource-constrained edge devices for real-time safety applications.
-MSTN achieves consistent SOTA performance across fundamentally different domains from behavioral analytics to mechanical systems, without any architectural modifications.
+MSTN-BiLSTM also shows strong performance in animal welfare evaluation [DISSANAYAKE2025111462], assessment with 93.00% accuracy on calf behavior recognition.
+Furthermore, MSTN-Transformer demonstrates better capability in mechanical prognostics, achieving 11.26 RMSE on the NASA Turbofan data set [NasaTurboFan] and significantly outperforming all baseline predictors in the estimate of the remaining useful life.
+[…]
 2.6 Ablation Study
-Table 4 presents an ablation study evaluating the effectiveness of the proposed MSTN-Transformer components across six datasets (ECL, Weather, ETTm1, ETTh2, Traffic, and ILI) and the MSTN-BiLSTM architecture components on the Weather, ETTm2, and Traffic datasets.
-To verify the contribution of each component, we compare the full MSTN-Transformer model against five variants: w/o CNN, which removes the convolutional pathway; w/o Transformer, which eliminates the core Transformer module; w/o SE, which removes the squeeze-and-excitation blocks; w/o MHTA, which excludes the Multi-Head Temporal Attention; and w/o Gated F., which removes the gated fusion mechanism.
-In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer to assess its importance in temporal sequence modeling.
-These results validate our architectural design choices and demonstrate the complementary nature of the integrated components.
-Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$.
-Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHTA w/o Gated F.
+Table 4: Ablation study of MSTN-Transformer and MSTN-BiLSTM. $H\in\{96,192,336,720\}$ and for ILI $H\in\{24,36,48,60\}$. Red and blue indicate first and second best results, respectively.
+Methods MSTN Full w/o CNN w/o Transf. w/o SE w/o MHA w/o SGF
 Metric MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE MSE MAE
@@
 96 0.041 0.161 0.045 0.165 0.048 0.173 0.042 0.163 0.043 0.164 0.045 0.168
-192 0.043 0.165 0.043 0.169 0.048 0.173 0.043 0.164 0.044 0.166 0.047 0.170
+192 0.042 0.163 0.046 0.169 0.048 0.173 0.045 0.166 0.044 0.166 0.047 0.170
 336 0.042 0.162 0.046 0.168 0.050 0.178 0.046 0.167 0.045 0.169 0.049 0.168
-720 0.047 0.173 0.051 0.179 0.055 0.186 0.049 0.172 0.050 0.178 0.053 0.179
+720 0.047 0.173 0.051 0.179 0.055 0.186 0.049 0.175 0.050 0.178 0.053 0.179
 Weather
-96 0.109 0.263 0.111 0.265 0.118 0.273 0.109 0.263 0.110 0.264 0.110 0.264
-192 0.111 0.264 0.115 0.269 0.120 0.275 0.111 0.264 0.111 0.265 0.112 0.265
-336 0.114 0.268 0.120 0.274 0.123 0.279 0.113 0.268 0.114 0.269 0.114 0.269
+96 0.109 0.263 0.112 0.266 0.118 0.273 0.110 0.264 0.111 0.265 0.112 0.265
+192 0.111 0.264 0.115 0.269 0.120 0.275 0.112 0.265 0.113 0.265 0.112 0.265
+336 0.114 0.268 0.120 0.274 0.123 0.279 0.115 0.269 0.116 0.270 0.116 0.270
 720 0.128 0.286 0.138 0.296 0.136 0.294 0.130 0.287 0.131 0.289 0.130 0.287
 ETTm1
-96 0.052 0.174 0.055 0.175 0.089 0.227 0.061 0.185 0.058 0.178 0.055 0.175
-192 0.103 0.242 0.105 0.243 0.105 0.243 0.109 0.246 0.098 0.233 0.096 0.229
-336 0.127 0.271 0.129 0.272 0.156 0.307 0.140 0.283 0.151 0.296 0.147 0.290
+96 0.052 0.174 0.062 0.186 0.089 0.227 0.061 0.185 0.057 0.178 0.055 0.177
+192 0.104 0.242 0.106 0.244 0.105 0.243 0.109 0.247 0.103 0.244 0.105 0.245
+336 0.127 0.271 0.146 0.289 0.156 0.307 0.140 0.283 0.151 0.296 0.147 0.290
 720 0.151 0.301 0.155 0.306 0.172 0.319 0.142 0.288 0.159 0.308 0.158 0.307
 ETTh2
-96 0.297 0.429 0.370 0.492 0.365 0.485 0.374 0.482 0.377 0.491 0.390 0.485
-192 0.300 0.436 0.404 0.511 0.391 0.505 0.445 0.529 0.382 0.498 0.395 0.501
-336 0.377 0.495 0.415 0.520 0.388 0.501 0.396 0.504 0.387 0.498 0.396 0.505
-720 0.400 0.509 0.427 0.522 0.384 0.498 0.426 0.523 0.426 0.531 0.439 0.531
+96 0.297 0.429 0.370 0.492 0.385 0.498 0.374 0.482 0.377 0.491 0.390 0.485
+192 0.300 0.436 0.404 0.511 0.391 0.505 0.392 0.506 0.382 0.496 0.395 0.501
+336 0.377 0.495 0.415 0.520 0.410 0.518 0.396 0.508 0.387 0.499 0.396 0.505
+720 0.400 0.509 0.428 0.524 0.430 0.525 0.426 0.523 0.426 0.531 0.439 0.531
 Traffic
 96 0.017 0.110 0.020 0.113 0.026 0.128 0.019 0.112 0.021 0.116 0.020 0.112
-192 0.018 0.111 0.020 0.115 0.026 0.127 0.020 0.113 0.023 0.116 0.021 0.116
-336 0.020 0.112 0.022 0.118 0.029 0.135 0.021 0.115 0.024 0.118 0.023 0.114
+192 0.018 0.111 0.020 0.115 0.026 0.127 0.020 0.113 0.023 0.116 0.020 0.113
+336 0.020 0.112 0.022 0.118 0.029 0.135 0.021 0.115 0.024 0.118 0.021 0.114
 720 0.020 0.113 0.026 0.128 0.026 0.128 0.022 0.118 0.025 0.119 0.025 0.120
 ILI
-24 0.129 0.254 0.140 0.272 0.139 0.274 0.132 0.256 0.131 0.253 0.133 0.256
+24 0.129 0.254 0.140 0.272 0.139 0.274 0.131 0.255 0.131 0.255 0.133 0.256
 36 0.155 0.287 0.167 0.301 0.168 0.307 0.161 0.298 0.157 0.293 0.168 0.304
-48 0.180 0.323 0.208 0.344 0.193 0.332 0.187 0.327 0.183 0.325 0.191 0.328
-60 0.190 0.332 0.197 0.338 0.199 0.339 0.193 0.333 0.193 0.332 0.194 0.338
+48 0.180 0.323 0.208 0.344 0.193 0.332 0.182 0.324 0.183 0.325 0.191 0.328
+60 0.190 0.332 0.197 0.338 0.199 0.339 0.193 0.333 0.194 0.334 0.194 0.338
 MSTN BiLSTM
-Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHTA w/o Gated F.
+Methods MSTN Full w/o CNN w/o BiLSTM w/o SE w/o MHA w/o SGF
 Weather
-96 0.110 0.264 0.112 0.266 0.119 0.274 0.110 0.264 0.110 0.263 0.110 0.264
-192 0.111 0.265 0.1132 0.267 0.120 0.275 0.112 0.265 0.111 0.265 0.112 0.266
-336 0.115 0.269 0.116 0.270 0.123 0.279 0.115 0.269 0.115 0.269 0.115 0.269
-720 0.132 0.289 0.138 0.292 0.139 0.296 0.132 0.293 0.133 0.295 0.134 0.298
+96 0.110 0.264 0.112 0.266 0.119 0.274 0.111 0.264 0.111 0.265 0.112 0.266
+192 0.111 0.265 0.113 0.267 0.120 0.275 0.112 0.266 0.113 0.267 0.113 0.266
+336 0.115 0.269 0.116 0.270 0.123 0.279 0.116 0.270 0.117 0.271 0.118 0.272
+720 0.132 0.289 0.138 0.295 0.139 0.296 0.134 0.294 0.135 0.294 0.134 0.298
 ETTm2
-96 0.238 0.390 0.243 0.400 0.264 0.430 0.262 0.394 0.240 0.388 0.261 0.399
-192 0.426 0.514 0.501 0.560 0.528 0.617 0.436 0.540 0.419 0.535 0.411 0.512
+96 0.238 0.390 0.243 0.400 0.264 0.430 0.262 0.394 0.240 0.392 0.261 0.399
+192 0.426 0.514 0.501 0.560 0.528 0.617 0.408 0.510 0.419 0.535 0.411 0.512
 336 0.517 0.594 0.581 0.623 0.571 0.612 0.535 0.592 0.536 0.593 0.626 0.655
-720 0.576 0.614 0.592 0.635 0.578 0.619 0.681 0.661 0.829 0.738 0.641 0.647
+720 0.576 0.614 0.660 0.651 0.671 0.665 0.638 0.645 0.829 0.738 0.641 0.647
 Traffic
-96 0.085 0.202 0.086 0.202 0.102 0.222 0.085 0.202 0.087 0.204 0.087 0.205
-192 0.087 0.205 0.088 0.207 0.101 0.221 0.087 0.205 0.088 0.208 0.088 0.209
-336 0.086 0.204 0.086 0.205 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
-720 0.087 0.207 0.089 0.210 0.102 0.224 0.092 0.218 0.091 0.217 0.093 0.218
+96 0.085 0.202 0.089 0.208 0.090 0.209 0.086 0.203 0.087 0.204 0.087 0.205
+192 0.087 0.205 0.092 0.218 0.101 0.221 0.088 0.207 0.088 0.208 0.088 0.209
+336 0.086 0.204 0.095 0.224 0.105 0.227 0.089 0.213 0.090 0.214 0.090 0.215
+720 0.087 0.207 0.098 0.227 0.107 0.229 0.091 0.218 0.091 0.217 0.093 0.218
+An ablation study systematically removes individual components from the full model to evaluate their contributions.
+Table 4 presents the results for MSTN-Transformer across six diverse benchmarks ECL (electricity), Weather, ETTm1, ETTh2, Traffic, ILI and for MSTN-BiLSTM on three datasets: Weather, ETTm2, and Traffic.
+For MSTN-Transformer, five variants are tested: w/o CNN, w/o Transformer, w/o SE, w/o MHA, and w/o SGF.
+While In MSTN-BiLSTM architecture, w/o BiLSTM denotes the variant of MSTN-BiLSTM without the BiLSTM layer.
+In the MSTN-Transformer variant, the results consistently show that the full MSTN-Transformer model achieves the best performance (ranked first in MSE/MAE) across nearly all prediction horizons and datasets.
+Removing the core Transformer module (w/o Transformer) leads to the most severe performance drop in most cases.
+For instance, on ETTh2 (a challenging dataset with high volatility), the MSE increases substantially from 0.297 (full model) to 0.385 at horizon 96 when the Transformer is ablated, highlighting its critical role in modeling complex temporal patterns.
+Similarly, the CNN pathway (w/o CNN) is highly important, especially on datasets like Traffic and ECL where local features are crucial; its removal causes noticeable degradation (e.g., Traffic H=96 MSE rises from 0.017 to 0.020).
+The SE blocks (w/o SE) also contribute significantly, often yielding the second-best or competitive results when present.
+For example, on Weather, the SE-ablated variant remains close to the full model, but on ETTm1, its removal increases MSE from 0.127 to 0.140 at H=336.
+The MHA (w/o MHA) and self-gated fusion (w/o SGF) modules further enhance robustness, with their absence leading to consistent, though sometimes smaller, declines across datasets.
+For MSTN-BiLSTM, the full model outperforms all ablated variants.
+Removing the BiLSTM layer (w/o BiLSTM) results in significant deterioration, particularly on complex datasets like ETTm2 where MSE at H=720 jumps from 0.576 to 0.671.
+This underscores the importance of recurrent modeling in capturing sequential dependencies.
+In summary, the ablation study validates that each architectural component is essential.
+The Transformer or BiLSTM, convolutional pathway, and attention-based mechanisms operate synergistically, with the full integrated model delivering the best overall forecasting accuracy across diverse temporal datasets.
 2.7 Model Complexity and Edge-AI Deployability
-The MSTN-Transformer and MSTN-BiLSTM architectures exemplify a fundamental design trade-off in temporal modeling.
-Both architectures integrate efficient parallel components CNN pathways ($\mathcal{O}(T)$) with either Transformer ($\mathcal{O}(T^{2})$) or BiLSTM ($\mathcal{O}(T)$) mechanisms, combined with $\mathcal{O}(T)$ fusion operations and finalized by a multi-head temporal attention layer of $\mathcal{O}(T^{2})$ complexity.
-This quadratic component dictates the overall asymptotic complexity of $\mathcal{O}(T^{2})$.
-However, this design preserves robust temporal modeling capabilities and achieves improved performance and practical efficiency compared to pure $\mathcal{O}(T^{2})$ Transformer-based models.
-Linear models such as DLinear and TS-Mixer achieve optimal $\mathcal{O}(T)$ complexity, while spectral and frequency-based approaches such as FEDformer and TimesNet attain $\mathcal{O}(T)$ and $\mathcal{O}(T\log T)$ complexity, respectively.
-The recently proposed MTST further reduces the cost to $\mathcal{O}(J_{b_{n}^{\prime}}^{2})$ via patching.
-Although efficient, these methods often trade off capacity for unconstrained long-range dependency modeling.
-Following the results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our results with their metrics reported in Table 5.
-The MSTN-Transformer achieves improved performance with 0.020 MSE and 0.72 ms inference time, representing 22.4× accuracy improvement and 34.7× speedup over SOFTS, achieving a favorable balance of accuracy and speed.
-This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
-Table 5: A: Performance Analysis (Traffic Dataset, $L=96$, $H=720$). B: MSTN-Trans. Inference Time Component Contributions. Red/Blue: First/Second ranks.
-MSE Perf. Gap Inf. Time (ms) Speed Gap
-A: SOTA MSE and Efficiency
+2.7.1 Complexity Analysis
+In this analysis, we denote the length of the lookback window as $L$, the number of variate features (channels) as $C$, and the task-specific output dimension as $K$.
+We use $C$ for channels to align with notation in common literature such as SOFTS [han2024softsefficientmultivariatetime], where $K$ is often referred to as the forecasting horizon ($H$).
+We simplify the asymptotic complexity expressions by omitting the model dimension $d$, assuming $d\approx C$.
+The MSTN framework addresses the fundamental trade-off in temporal modeling between representational capacity and computational efficiency.
+Although the proposed variants exhibit distinct encoding complexities with MSTN-BiLSTM maintaining strict linearity with respect to sequence length and MSTN-Transformer incurring a quadratic cost.
+They both converge on a unified efficiency strategy: ETA.
+This is achieved because both the parallel CNN and sequence modeling pathways utilize immediate temporal pooling to eliminate sequence-length dependence in subsequent layers.
+Specifically, both the parallel CNN and sequence modeling pathways (Transformer/BiLSTM) independently condense their temporal features via Global Average Pooling and Sequence Mean Pooling, respectively (reducing the sequence dimension from $L$ to $1$).
+Following concatenation, the subsequent SGF, SE-Block, and MHA refinement stages operate on static feature vectors.
+This strategic design ensures that the refinement pipeline scales as $\mathcal{O}(C^{2})$ with respect to channels but remains constant-time ($\mathcal{O}(1)$) with respect to the input length $L$, effectively isolating the heavy temporal computation to the initial stage.
+Component-Wise Analysis of MSTN-Transformer: The complexity is dominated by the sequence encoding stage.
+First, the Multi-Scale CNN extracts local features with linear complexity $\mathcal{O}(CL)$.
+Second, the Transformer Encoder applies self-attention; as this requires computing an $L\times L$ attention matrix, it introduces the dominant quadratic term $\mathcal{O}(CL^{2})$.
+Following this, Temporal Pooling aggregates the sequence with linear cost $\mathcal{O}(CL)$.
+Crucially, the Refinement stage operates on the pooled vector, scaling with channel interactions as $\mathcal{O}(C^{2})$.
+Finally, the Projection Head maps features to the output with complexity $\mathcal{O}(CK)$.
+Summing these components, the total asymptotic complexity is $\mathcal{O}(CL^{2}+C^{2}+CK)$, which scales quadratically with sequence length $L$.
+Component-Wise Analysis of MSTN-BiLSTM: In contrast, the MSTN-BiLSTM variant achieves strict linear efficiency with respect to sequence length.
+The BiLSTM Encoder processes the sequence recurrently without generating a global attention map, resulting in a linear complexity of $\mathcal{O}(CL)$.
+Similar to the Transformer variant, Temporal Pooling ($\mathcal{O}(CL)$) condenses the time dimension.
+The Refinement stage contributes $\mathcal{O}(C^{2})$, and the Projection Head contributes $\mathcal{O}(CK)$.
+Since the encoder avoids quadratic operations in $L$, the total start-to-end complexity remains highly efficient at $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with $L$.
+Table 5: Comprehensive evaluation of MSTN across diverse time-series benchmarks.
+Inference time of MSTN compared to models including SOFTS, iTransformer, TimesNet and PatchTST.
+Red/Blue: First/Second ranks.
+(a) Comparison of MSTN with the MSE and inference time (I.T) metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic dataset ($L=96$, $H=720$).
+MSE Perf. Gap I.T.(ms) Speed Gap
 […]
@@
 […]
-B: MSTN-Trans. Component Contributions Ablation Study (Traffic Dataset, $H=720$)
-Variant Size (MB) Time (ms) Speed Gap Role
+(b) Comprehensive evaluation of MSTN across diverse time-series forecasting benchmarks [wu2022autoformer].
+Dataset Horizon Size (MB) I.T. (ms)
+Traffic 96,192,336,720 2.87 0.72
+ETTh2 96,192,336,720 3.65 0.91
+ETTm2 96,192,336,720 2.92 0.85
+Weather 96,192,336,720 2.85 0.70
+Exchange 96,192,336,720 2.96 0.80
+ILI 24,36,48,60 2.51 0.63
+(c) Inference time (I.T., in ms) of MSTN variants compared to TimesNet and PatchTST across seven international datasets.
+Dataset MSTN Size (MB) I.T. TimesNet I.T. PatchTST I.T.
+Rodegast [DARUS33012023] MSTN-Trans. 0.54 0.15 5.07 4.80
+MSTN-BiLSTM 4.39 0.81
+Boubezoul [BOUBEZOUL2020105577] MSTN-Trans. 0.96 0.24 4.20 4.68
+MSTN-BiLSTM 4.16 2.79
+UCI-HAR [har240] MSTN-Trans. 29.98 3.48 45.30 4.36
+MSTN-BiLSTM 7.14 2.03
+PAMAP2 [pamap2_physical_activity_monitoring_231] MSTN-Trans. 1.58 0.20 0.46 3.21
+MSTN-BiLSTM 0.76 0.086
+ActBeCalf [DISSANAYAKE2025111462] MSTN-Trans. 0.89 0.14 8.53 2.20
+MSTN-BiLSTM 4.14 1.76
+MetroPT3 [metropt3] MSTN-Trans. 1.30 0.19 0.09 0.19
+MSTN-BiLSTM 0.57 0.042
+NASA [NasaTurboFan] MSTN-Trans. 1.31 0.45 4.52 2.31
+MSTN-BiLSTM 4.25 1.48
+2.7.2 Comparative Analysis and Edge-AI Deployability
+Recent efficient architectures demonstrate varying complexity profiles: linear models such as DLinear [zeng2022transformerseffectivetimeseries] achieve $\mathcal{O}(CL)$ complexity, while spectral approaches like FEDformer [zhou2022fedformer] and TimesNet [wu2023timesnettemporal2dvariationmodeling] achieve $\mathcal{O}(CL\log L)$ complexity.
+Following the notation in SOFTS [han2024softsefficientmultivariatetime], their model achieves $\mathcal{O}(CL+CH)$ complexity via stochastic pooling, while iTransformer ($\mathcal{O}(C^{2}+CL+CH)$), PatchTST ($\mathcal{O}(CL^{2}+CH)$), and Standard Transformers ($\mathcal{O}(N\cdot CL^{2})$) incur heavier costs.
+Against this landscape, the proposed MSTN-BiLSTM achieves strict linear temporal complexity $\mathcal{O}(CL+C^{2}+CK)$.
+Theoretically, this matches the optimal efficiency of SOFTS in the predominant real-world scenario where sequence length exceeds channel dimension ($L\gg C$).
+While SOFTS retains a theoretical advantage in high-dimensional, short-sequence settings ($C\gg L$) due to its purely linear channel dependence, such data structures are rare in long-term time series forecasting tasks.
+Meanwhile, MSTN-Transformer scaling as $\mathcal{O}(CL^{2}+C^{2}+CK)$ optimizes practical runtime by confining the quadratic term to a single encoder layer rather than a deep stack ($N$ layers) as seen in iTransformer or standard Transformer architectures.
+This strategic design allows MSTN to balance the high-capacity modeling of Transformers with the latency requirements of Edge-AI, delivering state-of-the-art performance with minimized computational overhead.
+One thing to note is that despite the MSTN-BiLSTM’s better theoretical scaling ($\mathcal{O}(L)$) than the MSTN-Transformer’s $\mathcal{O}(L^{2})$, the latter runs significantly faster than the former (Table 5).
+This paradox is due to the fundamental conflict between the model’s structure and the way modern acceleration architectures process the data.
+The BiLSTM encoder is inherently recurrent, meaning that the calculation of the hidden state $h_{t}$ is strictly dependent on the previous state $h_{t-1}$.
+This sequential data dependency prevents the hardware’s parallel processing units from being fully utilized, forcing the computation into a slow, iterative loop (low hardware utilization).
+In contrast, the Transformer executes its $\mathcal{O}(L^{2})$ self-attention via highly optimized [vaswani2023attentionneed], complete matrix parallelization.
+The core operation, the $QK^{T}$ matrix multiplication, is performed simultaneously across thousands of processing elements.
+This high utilization of the hardware’s parallel capacity, confined to a single encoder layer before pooling, ensures the Transformer’s execution speed drastically outweighs the theoretical linearity of the BiLSTM’s sequential cost, resulting in lower practical latency across high-throughput platforms.
+Standard Transformers maintain the full sequence length $L$ across a multi-layer stack of $N$ blocks for self-attention operations, resulting in cumulative attention cost proportional to $N\cdot\mathcal{O}(CL^{2})$.
+In contrast, MSTN applies attention in its encoder, then immediately pools ($L\to 1$).
+Standard Transformers maintain $L$ across all layers, while MSTN eliminates $L$-dependence after encoding.
+Consequently, the remainder of the network operates on static feature vectors with $\mathcal{O}(1)$ complexity relative to $L$ for operations involving sequence length.
+By eliminating the recursive application of attention across deep layers, MSTN reduces the sequence-dependent FLOPs significantly.
+While MSTN-BiLSTM avoids quadratic attention costs entirely by employing BiLSTM encoding with $\mathcal{O}(CL)$ complexity.
+When combined with the remaining pipeline components (CNN, refinement, and projection), the total complexity is $\mathcal{O}(CL+C^{2}+CK)$, which scales linearly with sequence length $L$.
+Following the baseline results and inference time metrics reported in SOFTS [han2024softsefficientmultivariatetime] on the Traffic data set (lookback window $L=96$, horizon $H=720$), we compare our MSTN results with their metrics reported in Table 5.
+MSTN-Transformer demonstrates that architectural efficiency can surpass asymptotic predictions.
+MSTN-Transformer achieves $69\times$ faster inference than PatchTST (0.72 ms vs.
+50.00 ms).
+Furthermore, it outperforms models with better theoretical scaling, showing speedups of $34.7\times$ over SOFTS and $305\times$ over TimesNet.
+This practical gain, irrespective of the asymptotic class, is primarily delivered by the early temporal pooling of the MSTN architecture ($L\to 1$), a strategic design choice that dramatically reduces the overall FLOPs compared to deep, unpooled attention stacks.
+This performance represents a $22.4\times$ accuracy improvement and $34.7\times$ speedup over the SOFTS baseline, and a significant speed and accuracy advantage over all other models detailed in Table 5.
+This improved efficiency enables substantially reduced memory usage and inference latency, positioning MSTN as a practical solution for deployment on constrained hardware platforms.
+As summarized in Table 5, MSTN demonstrates consistently efficient inference across six widely used long-term forecasting benchmarks.
+Despite the variation in dataset scale and temporal structure, MSTN maintains a compact model size in the range of 2.5–3.7 MB and sustains sub-millisecond latency across all horizons, including the longer 336 and 720-step settings.
+This stability in runtime behavior across heterogeneous datasets such as Traffic, ETTh2, ETTm2, Weather, Exchange, and ILI highlights the robustness of MSTN’s dual-path, multi-scale feature extraction.
+The uniformly low inference cost emphasizes its suitability for real-time forecasting scenarios and deployment in resource-constrained environments where computational consistency is critical.
 […]
+3 Discussion
+This work introduces MSTN, a DL framework founded on a hierarchical multi-scale modeling principle and defined by its strategic use of ETA.
+MSTN integrates core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; a sequential modeling component for long-term global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a SGF mechanism augmented with SE and MHA to enable dynamic, context-aware feature integration.
+Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures).
+These limitations have historically forced a compromise between scale diversity and inference speed.
+In contrast, our approach efficiently integrates convolutional local pattern extraction with long-term global dependency modeling, leveraging ETA via an adaptive SGF mechanism.
+This enables our model to capture multi-scale temporal features while maintaining computational efficiency and simultaneously achieving SOTA performance.
 […]
-[…]
-[…]
-[…]
-[…]
-As shown in Table 3, MSTN-Transformer delivers competitive performance across domains, attaining 99.53% accuracy on Rodegast [DARUS33012023] while maintaining efficient inference at 0.155 ms per sample with only 0.54 MB model size.
-This represents up to 33× speedup over TimesNet (5.07 ms) and 31.0× over PatchTST (4.80 ms), demonstrating favorable performance-efficiency trade-offs across diverse datasets.
-The combination of compact model sizes (0.54–7.14 MB), improved accuracy, and millisecond-level latency positions MSTN as an ideal solution for complex multivariate time series analysis tasks in edge computing environments.
-These results demonstrate that capable temporal models can achieve practical deployability in edge environments.
-The small model footprint satisfies memory constraints, millisecond-level latency enables real-time operation, and high accuracy ensures reliability.
-This collectively addresses core edge-AI challenges for applications like safety monitoring, predictive maintenance, and activity recognition.
-3 Discussion
-This work introduces MSTN, a DL architecture founded on a hierarchical multi-scale modeling principle.
-It integrates three core components: a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; sequential modeling component for global dependency modeling, empirically validated with BiLSTM and Transformer variants; and a gated fusion mechanism augmented with SE and MHTA to enable dynamic, context-aware feature integration.
-Existing methods often rely on fixed-scale structural priors (e.g., patch-based tokenization, fixed frequency transformations, or frozen backbone architectures), forcing a compromise between scale diversity and inference speed.
-In contrast, our approach efficiently integrates convolutional local pattern extraction with global dependency modeling via an adaptive gated fusion mechanism.
-This enables our model to capture multi-scale temporal features without compromising computational efficiency, ultimately achieving state-of-the-art performance.
-Comprehensive evaluations demonstrate that the MSTN-Transformer establishes a new SOTA across forecasting, imputation, classification, and generalizability.
-It achieves a 22.4× improvement in accuracy over the SOFTS model while being 34.7× faster in inference, all within a compact 2.87 MB footprint, while MSTN BiLSTM shows competitive performance.
-This dual advancement in both performance and efficiency is a key differentiator from prior work.
-The model also exhibits exceptional cross-domain generalization, with its small model size (0.54–7.14 MB) making it suitable for edge deployment.
-Ablation studies confirm that each component is critical, with the transformer core providing a foundational capability and the multi-scale and fusion modules delivering significant competitive gains.
-Future work will explore cross-variate attention mechanisms, and large-scale pre-training to further enhance MSTN’s capabilities as a foundation model for time series understanding.
-Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time forecasting applications across diverse domains from healthcare to industrial monitoring.
+Although MSTN excels in standard forecasting regimes ($L\gg C$), we acknowledge that factorization-based models such as SOFTS retain a theoretical efficiency advantage in high-dimensional, short-sequence settings ($C\gg L$).
+However, since such datasets are rarely encountered in long-term time series forecasting tasks, this limitation has a negligible impact on practical deployability.
+Future work will explore cross-variate attention mechanisms and large-scale pre-training to further enhance the capabilities of MSTN as a foundation model for time series understanding.
+Our work paves the way for efficient multi-scale temporal modeling under resource constraints, enabling real-time applications across diverse domains from healthcare to industrial monitoring.
 4 Methods
+This section details the MSTN architecture and its methodological innovations, followed by the baseline models, and training configurations.
 4.1 Proposed Architecture: MSTN
@@
 Fig. 1: Proposed MSTN: (a) architectural diagram and (b) signal processing pipeline
-The proposed MSTN (Fig.
-1a) is a novel hybrid DL architecture founded on a hierarchical multi-scale modeling principle.
-It uses a dual-branch processing pipeline: (i) a convolutional branch that extracts detailed local motion patterns, and (ii) sequence modeling core that learns temporal dependencies in the time-series data, which we empirically validate using both BiLSTM and Transformer variant in forecasting, imputation, classification and generalizability studies.
-(iii) The outputs of both branches are integrated using a learnable gated fusion mechanism, followed by channel-wise recalibration through SE blocks and MHTA, enabling the model to focus on critical, risk-indicating time steps.
-The fused representation is normalized, regularized with dropout, and linearly projected to produce probabilities.
-Training takes advantage of focal loss to mitigate class imbalance.
-Fig.
-1b shows the processing pipeline of time series signals.
-CNN pathway captures local patterns, while BilSTM and Transformer captures long-range temporal dependencies in the time-series data, with gated fusion adaptively combining both feature streams.
-4.1.2 Model Structure
-As aforementioned, based on the multi-scale nature of time series, we propose MSTN with a parallel architecture to capture temporal patterns at different resolutions.
-For comprehensive temporal modeling, we design dual pathways within MSTN that process the time series through complementary mechanisms and simultaneously model both local and global variations through a parameter-efficient novel hybrid block.
-The MSTN architecture and signal processing pipeline are illustrated in Fig.
-1.
-We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times M}$ with lookback window sequence length $L$ and $M$ variates features, MSTN learns rich temporal representations for multiple tasks including forecasting, imputation and classification through a novel parallel architecture.
+The proposed Multi-Scale Temporal Network (MSTN; Fig.
+1a) is a hybrid deep-learning architecture designed around a parallel multi-scale modeling paradigm.
+MSTN processes temporal signals through two coordinated encoding pathways, each specialized for complementary aspects of temporal structure.
+A key design element is the use of an early temporal aggregation (ETA) mechanism, applied directly to the outputs of the dual encoders, which collapses the full input sequence of length $L$ into a single fixed-dimensional representation ($L\rightarrow 1$).
+This operation ensures that all subsequent refinement stages operate in constant time with respect to sequence length, thereby enabling efficient processing of long temporal inputs.
+The network comprises three principal modules.
+(i) A multi-scale convolutional branch captures localized temporal patterns using hierarchically stacked convolutions; its output is aggregated via global average pooling, promoting robustness and translation invariance.
+(ii) A sequence modeling branch captures long-range dependencies, instantiated using either a bidirectional LSTM (MSTN-BiLSTM) or a Transformer encoder (MSTN-Transformer).
+To maintain computational efficiency, this branch is compressed through sequence mean pooling, yielding a global summary vector.
+(iii) A multi-scale fusion and refinement module integrates the pooled representations from the two branches.
+The features are concatenated and passed through a scale-gated fusion (SGF) mechanism, which adaptively modulates the contributions of each pathway.
+The fused representation is subsequently refined using channel-wise recalibration via squeeze–excitation blocks and multi-head attention, enabling the network to emphasize the most informative feature dimensions.
+The final representation is normalized, regularized through dropout, and projected to the task-specific output space.
+To address class imbalance, training is performed using a focal loss objective.
+Figure 1b provides an overview of the end-to-end processing pipeline, from parallel feature extraction to multi-scale fusion and refinement.
+4.1.2 Methodology
+We consider the following problem: Given multivariate time series samples $X_{1:L}\in\mathbb{R}^{L\times C}$ with a lookback window sequence length $L$ and $C$ variate features, MSTN learns temporal representations through a novel parallel architecture.
+We detail the complete process below.
 Forward Process.
-The input $X\in\mathbb{R}^{B\times T\times D}$ is processed through dual complementary pathways that operate simultaneously.
-The CNN pathway captures local temporal patterns while the Transformer pathway models global dependencies or BiLSTM models sequential intelligence, with their outputs fused through sophisticated gated mechanisms for enhanced representation learning.
-To capture multi-scale temporal variations in length-$T$ time series with $D$ variates organized as $X_{1D}\in\mathbb{R}^{B\times T\times D}$, the architecture processes inputs through two parallel pathways with convolutional pathway and sequence modeling (Transformer or BiLSTM pathway).
-The convolutional pathway captures fine-grained, short-term temporal patterns through hierarchical feature learning.
-The input undergoes successive 1D convolutions to extract multi-scale features, expanding the feature dimension while preserving temporal resolution, followed by refinement of the representations.
+The input $X\in\mathbb{R}^{B\times L\times C}$ is processed through dual complementary pathways that operate simultaneously.
+The CNN pathway captures local temporal patterns, while the sequence modeling pathway (implemented as either a Transformer or a BiLSTM) captures long-range temporal dependencies.
+Their outputs are fused through a sophisticated gating and attention mechanism for enhanced representation learning.
+The input undergoes successive 1D convolutions to extract multi-scale features, adjusting the feature dimension while preserving temporal resolution, followed by refinement of the representations.
+The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training.
 To aggregate these temporal features into a compact representation, we apply global average pooling across the entire sequence:
-$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{Conv1D}_{7}(X_{1D}))\in\mathbb{R}^{B\times 128\times T}$ (1)
-$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times T}$ (2)
+$H_{\text{conv}}^{(1)}$ $=\text{ReLU}(\text{BN}(\text{Conv1D}_{7}(X^{T})))\in\mathbb{R}^{B\times 128\times L}$ (1)
+$H_{\text{conv}}^{(2)}$ $=\text{ReLU}(\text{Conv1D}_{5}(H_{\text{conv}}^{(1)}))\in\mathbb{R}^{B\times 64\times L}$ (2)
 $\mathbf{z}_{\text{cnn}}$ $=\text{GlobalAvgPool1D}(H_{\text{conv}}^{(2)})\in\mathbb{R}^{B\times 64}$ (3)
 where each of the 64 feature channels is summarized by its average activation over time through temporal pooling.
-The $\text{Conv1D}_{k}$ convolutional layers employ a hierarchical design where the first layer with kernel size $k=7$ extracts basic temporal patterns, and the second layer with kernel size $k=5$ learns combinations of these patterns into more complex features, with padding to maintain temporal dimensions, and batch normalization stabilizes training.
-We empirically validate two complementary sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM.
-Features from both pathways are elegantly concatenated, uniting the CNN’s local precision with either Transformer or BiLSTM, to capture long-range temporal dependencies.
-When using Transformer: The architecture captures long-range dependencies through a multi-layer transformer encoder.
-The input is processed through self-attention mechanisms where each of the T time steps maintains a 64-dimensional representation enriched with global contextual information.
+We employ two sequence modeling approaches: MSTN-Transformer and MSTN-BiLSTM.
+In MSTN-Transformer, the architecture captures long-range dependencies through a single Transformer encoder block composed of 4 layers with 8 attention heads.
+The input is processed through self-attention mechanisms where each of the L time steps maintains a 128-dimensional representation enriched with global contextual information.
 To consolidate these temporal representations, we apply sequence mean pooling across the time dimension:
-$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times T\times 64};\quad\mathbf{z}_{\text{trans}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{trans},t}\in\mathbb{R}^{B\times 64}$ (4)
-The transformer encoder uses 4 layers with 8 attention heads.
-When using BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
-$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times T\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{T}\sum_{t=1}^{T}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
+$H_{\text{trans}}=\text{TransformerEnc}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{trans}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{trans},t}\in\mathbb{R}^{B\times 128}$ (4)
+In MSTN-BiLSTM: The architecture captures bidirectional sequential dependencies, processing sequences in both forward and backward directions to learn complex temporal dynamics and memorize long-range patterns:
+$H_{\text{bilstm}}=\text{BiLSTM}(X)\in\mathbb{R}^{B\times L\times 128};\quad\mathbf{z}_{\text{bilstm}}=\frac{1}{L}\sum_{t=1}^{L}H_{\text{bilstm},t}\in\mathbb{R}^{B\times 128}$ (5)
 The BiLSTM employs 2 layers with 64 hidden units per direction, capturing both causal and anti-causal temporal relationships.
-The CNN pathway utilizes global average pooling to capture local pattern summaries, while the Transformer and BiLSTM pathways employ sequence mean pooling to preserve global contextual information.
-The parallel features undergo sophisticated fusion through a multi-stage enhancement process that intelligently combines the strengths of both pathways.
+The CNN pathway utilizes global average pooling, and the sequence modeling pathways employ sequence mean pooling.
+These pooling operations, along with the subsequent reshaping, collectively form the ETA module, which collapses the temporal dimension $L$ to $1$.
+This is the critical step that ensures subsequent operations happen at an $\mathcal{O}(1)$ inference cost.
+The parallel features undergo sophisticated fusion through a multi-stage enhancement process that combines the strengths of both pathways.
 For MSTN-Transformer:
-$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans}}]\in\mathbb{R}^{128}$ (6)
-where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{64}$. For MSTN-BiLSTM:
-$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{bilstm}}]\in\mathbb{R}^{192}$ (7)
-where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{bilstm}}\in\mathbb{R}^{128}$.
-The architecture then performs adaptive weighting through an intelligent gated fusion mechanism, illustrated in Fig.
+$\mathbf{z}_{\text{concat}}=[\mathbf{z}_{\text{cnn}};\mathbf{z}_{\text{trans/bilstm}}]\in\mathbb{R}^{192}$ (6)
+where $\mathbf{z}_{\text{cnn}}\in\mathbb{R}^{64}$ and $\mathbf{z}_{\text{trans}}\in\mathbb{R}^{128}$.
+The architecture then performs adaptive weighting through an SGF mechanism, illustrated in Fig.
 1, that learns to dynamically balance feature contributions:
 $\mathbf{z}_{\text{fused}}=\mathbf{z}_{\text{concat}}\odot\sigma(W_{g}\mathbf{z}_{\text{concat}}+b_{g})$ (7)
-where for MSTN-Transformer: $W_{g}\in\mathbb{R}^{128\times 128}$, $b_{g}\in\mathbb{R}^{128}$ and for MSTN-BiLSTM: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$.
-This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal intelligence.
-The fused features are reshaped for sequence processing: $z_{\text{fused\_seq}}=z_{\text{fused}}\text{.unsqueeze}(1)\text{.repeat}(1,T,1)\in\mathbb{R}^{B\times T\times d}$ where $d=128$ for MSTN-Transformer and $d=192$ for MSTN-BiLSTM.
-Channel-wise attention then enhances features through separate SE blocks for each variant: For MSTN-Transformer: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{16\times 128}$, $W_{2}\in\mathbb{R}^{128\times 16}$ with reduction ratio 8.
-For MSTN-BiLSTM: $z_{\text{se}}=z_{\text{fused\_seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\frac{1}{T}\sum_{t=1}^{T}z_{\text{fused\_seq},t}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8.
+where for both variants: $W_{g}\in\mathbb{R}^{192\times 192}$, $b_{g}\in\mathbb{R}^{192}$.
+This mechanism learns to emphasize the most relevant features from each pathway, creating a harmonious blend of multi-scale temporal features.
+The fused features are reshaped for efficient single-token processing: $\mathbf{z}_{\text{seq}}=\mathbf{z}_{\text{fused}}\otimes\mathbf{1}\in\mathbb{R}^{B\times 1\times 192}$.
+Channel-wise attention then enhances features through an SE block.
+Given the temporal dimension is already $L=1$, the global pooling step within SE is trivially executed: $\mathbf{z}_{\text{se}}=\mathbf{z}_{\text{seq}}\odot\sigma(W_{2}\text{ReLU}(W_{1}\mathbf{z}_{\text{seq}}))$ where $W_{1}\in\mathbb{R}^{24\times 192}$, $W_{2}\in\mathbb{R}^{192\times 24}$ with reduction ratio 8.
 The SE mechanism amplifies informative channels while suppressing less useful ones.
-Temporal relationships are refined through multi-head attention that captures interdependencies across the enhanced features:
-$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(T^{-1}\textstyle\sum_{t=1}^{T}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$
-with 4 attention heads (dim=64).
-The attention output undergoes layer normalization for stabilization and dropout for regularization, producing refined features $\mathbf{z}_{\text{final}}$ for the final classification head.
-The parallel pathway design enables MSTN to achieve comprehensive temporal modeling where $\mathcal{F}_{\text{CNN}}$ captures local patterns and $\mathcal{F}_{\text{Transformer}}$ models global dependencies, with hierarchical fusion creating synergistic representations that surpass individual pathway capabilities.
-This architectural innovation establishes a new paradigm for multi-scale time series analysis.
-4.1.3 Method-Level Innovation
-Many recent time-series models including PatchTST [nie2023timeseriesworth64] (patch-based tokenization), TimesNet [wu2023timesnettemporal2dvariationmodeling] (fixed frequency transformations), and LLM-based approaches like LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] (frozen backbone architectures) often rely on rigid, fixed-scale structural priors.
-This architectural inflexibility forces a fundamental trade-off: models either compromise on capturing multi-scale temporal diversity or suffer from high inference latency, rendering them suboptimal for real-time, resource-constrained deployment.
-The proposed MSTN addresses this core challenge through a novel multi-scale hierarchical architecture designed to capture both local patterns and long-range temporal dependencies without structural compromises.
-The architecture integrates three key components: (i) a multi-scale convolutional encoder that constructs a hierarchical feature pyramid; (ii) a sequence temporal modeling core, which we empirically validate with both BiLSTM and Transformer variants, establishing a flexible foundation for future architectural advancements; and (iii) a gated fusion mechanism augmented with SE and MHTA for dynamic, context-aware feature integration.
-This hierarchical design enables seamless adaptation to diverse temporal modeling paradigms while maintaining architectural consistency.
-Our comprehensive experiments demonstrate that these design choices allow MSTN to achieve improved performance compared to SOTA baselines while supporting low-latency inference critical for edge deployment.
-4.2 Dataset and Baselines
+Feature dependencies are refined through MHA.
+Since the temporal dimension is condensed ($L=1$), the MHA block functions here as a sophisticated global feature recalibration layer rather than a temporal mixer, confirming $\mathcal{O}(1)$ complexity:
+$\mathbf{z}_{\text{final}}=\text{Dropout}\left(\text{LayerNorm}\left(\sum_{t=1}^{1}\text{MHA}(\mathbf{z}_{\text{se}})_{t}\right),p=0.3\right)\in\mathbb{R}^{B\times 192}$ (8)
+where 192 is the feature dimension for both MSTN-Transformer and MSTN-BiLSTM, with 4 attention heads (dim=48 per head).
+The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification).
+The refined features $\mathbf{z}_{\text{final}}$ are then passed to the final task-specific head (imputation, forecasting, or classification).
+The prediction head consists of a single linear layer that maps the $192$-dimensional aggregated feature vector to the task-specific output shape.
+Task-Specific Prediction and Loss Functions: We detail the prediction head ($\hat{Y}$ or $\hat{X}$) and the corresponding loss function $\mathcal{L}$ for each of the three tasks:
+Time Series Imputation: The head reconstructs the input sequence $\hat{X}_{1:L}\in\mathbb{R}^{B\times L\times C}$, and the loss is calculated as masked MSE only over the observed values ($\Omega$ is the mask matrix):
+$\hat{X}_{1:L}$ $=W_{i}\mathbf{z}_{\text{final}}+b_{i}\in\mathbb{R}^{B\times L\times C}$ (9)
+$\mathcal{L}_{\text{imputation}}$ $=\frac{1}{|\Omega|}\sum_{i=1}^{B}\sum_{t=1}^{L}\sum_{c=1}^{C}\Omega_{i,t,c}\cdot\left(\hat{X}_{i,t,c}-X_{i,t,c}\right)^{2}$ (10)
+Long-Term Forecasting: The head produces a forecast $\hat{Y}_{1:T}\in\mathbb{R}^{B\times T\times C}$ for a horizon $T$, and the objective is to minimize the MSE:
+$\hat{Y}_{1:T}$ $=W_{f}\mathbf{z}_{\text{final}}+b_{f}\in\mathbb{R}^{B\times T\times C}$ (11)
+$\mathcal{L}_{\text{forecast}}$ $=\frac{1}{2TC}\sum_{i=1}^{B}\sum_{t=1}^{T}\sum_{c=1}^{C}\left(\hat{Y}_{i,t,c}-Y_{i,t,c}\right)^{2}$ (12)
+Time Series Classification: The head computes class probabilities $P\in\mathbb{R}^{B\times K}$ over $K$ classes, and the loss is the canonical Cross-Entropy (CE) Loss:
+$P$ $=\text{Softmax}(W_{c}\mathbf{z}_{\text{final}}+b_{c})\in\mathbb{R}^{B\times K}$ (13)
+$\mathcal{L}_{\text{classify}}$ $=-\frac{1}{B}\sum_{i=1}^{B}\sum_{k=1}^{K}Y_{i,k}\log(P_{i,k})$ (14)
+where $Y$ is the one-hot encoded true label. This structure allows the MSTN framework to be universally applied to various multivariate time series tasks.
+4.2 Method Level Innovation
+Table 6: Analysis of model architectures by design characteristics.
+This table outlines how different state-of-the-art models align with the established Channel Interaction (CI, CD, CP) strategies based on the temporal and channel characteristics they model.
+MSTN is positioned as a hybrid design (CI+CD).
+(Asym.: Asymmetry, Lag.: Lagginess, Pol.: Polarity, Gw.: group-wise, Dyn.: Dynamism, Ms.: Multi-scale)
+Strategy Mechanism Characteristic Method
+Asym. Lag. Pol. Gw. Dyn. Ms.
+CI - - - - - - - PatchTST
+CI - - - - - - - CycleNet
+CI - - - - - - - DLinear
+CI - - - - - - - Timer
+CI - - - - - - - Chronos
+CI - - - - - - - LLM4TS
+CI - - - - - - - Time-LLM
+CI - - - - - - - RevIN
+CD CNN-based $\checkmark$ - - - - - Informer
+CD CNN-based $\checkmark$ - - - - - Autoformer
+CD CNN-based $\checkmark$ - - - - - FEDformer
+CD CNN-based $\checkmark$ - - - - - TimesNet
+CD MLP-based $\checkmark$ - - - - - TSMixer
+CD MLP-based $\checkmark$ - - - - - TTM
+CD Transformer-based $\checkmark$ - - - $\checkmark$ - iTransformer
+CD Transformer-based $\checkmark$ - - - $\checkmark$ - Crossformer
+CD Transformer-based $\checkmark$ $\checkmark$ - - - - VCformer
+CD Transformer-based $\checkmark$ $\checkmark$ - - $\checkmark$ - MOIRAI
+CD Transformer-based $\checkmark$ - - - $\checkmark$ - UniTS
+CD GNN-based - - - $\checkmark$ - - GTS
+CD GNN-based $\checkmark$ - - - - $\checkmark$ MSGNet
+CD GNN-based - $\checkmark$ - - - - FourierGNN
+CD GNN-based - $\checkmark$ - - - - FC-STGNN
+CD GNN-based $\checkmark$ - - - $\checkmark$ - TPGNN
+CD Linear Channel Mixing (Factorization) $\checkmark$ - - - - - SOFTS
+CD Others $\checkmark$ - - - $\checkmark$ - C-LoRA
+CP CNN-based $\checkmark$ - - - - - ModernTCN
+CP Transformer-based $\checkmark$ - - $\checkmark$ - - DUET
+CP Transformer-based $\checkmark$ - - - $\checkmark$ - MCformer
+CP Transformer-based $\checkmark$ - - $\checkmark$ $\checkmark$ - DGCformer
+CP Transformer-based - - - - $\checkmark$ - CM
+CP GNN-based $\checkmark$ - - - - - MTGNN
+CP GNN-based $\checkmark$ - $\checkmark$ - - - CrossGNN
+CP GNN-based $\checkmark$ - - $\checkmark$ - - WaveForM
+CP GNN-based - - - - $\checkmark$ - MTSF-DG
+CP GNN-based $\checkmark$ - - $\checkmark$ - - ReMo
+CP GNN-based $\checkmark$ - - $\checkmark$ $\checkmark$ $\checkmark$ Ada-MSHyper
+CP Others $\checkmark$ $\checkmark$ - - - - LIFT
+CP Others $\checkmark$ - - $\checkmark$ - - CCM
+CI + CD Parallel CNN (CI) and BiLSTM/Transformer(CD) based Dual Encoder + ETA $\checkmark$ - - - $\checkmark$ $\checkmark$ MSTN
+The established field of multivariate time–series forecasting is conventionally organized around three channel interaction strategies—channel–independent (CI), channel–dependent (CD), and channel–projected (CP) approaches [qiu2025comprehensivesurveydeeplearning]—as summarized in Table 6.
+These strategies specify how a model encodes and exploits inter-variable dependencies across the $C$ channels of a multivariate signal.
+CI.
+CI methods (e.g., DLinear, PatchTST) process each variable independently using lightweight temporal operators such as linear layers or multilayer perceptrons.
+Their advantage lies in linear channel cost $\mathcal{O}(C)$ and efficient temporal complexity $\mathcal{O}(L)$, though this comes at the expense of discarding cross-channel structure.
+CD.
+CD approaches (e.g., iTransformer) explicitly model inter-variable correlation through dense attention or graph-based mechanisms.
+While these models offer comprehensive feature coupling, they incur high computational cost—typically $\mathcal{O}(C^{2})$ or $\mathcal{O}(L^{2})$—which limits their scalability and latency characteristics.
+CP.
+CP models (e.g., MCformer, MTGNN) map the channel dimension into a low-rank latent space before applying interaction, thereby avoiding the full $\mathcal{O}(C^{2})$ cost of CD models.
+Although channel-efficient, these models often impose simplified temporal assumptions, reducing their ability to capture multi-scale dynamics.
+Despite the usefulness of this taxonomy, it does not fully resolve the central challenges in long-sequence time–series applications, where models must simultaneously capture short- and long-range temporal dependencies while ensuring low-latency inference.
+Consequently, there is a need for architectures capable of representing multi-scale temporal structure without incurring prohibitive computational overhead.
+Existing SOTA systems illustrate these trade-offs.
+CI-based PatchTST [nie2023timeseriesworth64] relies on fixed patch-based tokenization; LLM-derived approaches such as LLM4TS [10.1145/3719207] and TIME-LLM [jin2024timellmtimeseriesforecasting] depend on frozen backbone architectures; CD-based TimesNet [wu2023timesnettemporal2dvariationmodeling] employs fixed frequency priors; and CP-based MCformer [10533212] attains efficiency through partial channel blending.
+These designs incorporate rigid structural priors that limit their ability to capture diverse temporal scales without sacrificing computational efficiency.
+To address these limitations, we introduce MSTN, a hybrid architecture grounded in the Efficient Temporal Aggregation (ETA) principle.
+MSTN explicitly targets the multi-scale nature of long-sequence data while avoiding the $\mathcal{O}(L^{2})$ bottleneck typical of Transformer-based models, enabling $\mathcal{O}(1)$ inference with respect to sequence length.
+By combining the complementary strengths of CI and CD models with a novel temporal aggregation mechanism, MSTN delivers strong performance across forecasting, imputation, and classification tasks while maintaining the low-latency characteristics required for real-world deployment.
+4.2.1 Core Innovation: Dual-Path Design and ETA
+Table 7: Comparison of architectural characteristics.
+This table contrasts the MSTN framework against the three standard Channel Interaction strategies (CI, CD, CP) across critical performance dimensions.
+MSTN achieves high capacity while neutralizing the $\mathcal{O}(L^{2})$ computational bottleneck.
+Dimension CI CD CP (CI+CD) MSTN Rationale
+Efficiency High Low Moderate High ETA enables $\mathcal{O}(1)$ refinement after initial encoding.
+Robustness High Low Moderate High Dual-encoder with SGF provides complementary feature stability.
+Generaliza- bility Low Moderate High High Achieved via a single, aggregated feature vector ($\mathbf{z}_{\text{final}}$) that supports multi-task learning (Forecasting, Imputation, Classification).
+Capacity Low High Moderate High Transformer/BiLSTM path captures global dependencies.
+Ease of Implementation High Moderate Low Moderate Relies on standard components (CNN, Transformer/BiLSTM, SGF, SE, MHA) rather than complex, specialized graph or recursive structures.
+The MSTN framework integrates a dual-path temporal encoding architecture with the ETA mechanism, yielding a principled hybrid approach that reconciles the core trade-offs among CI-, CD-, and CP-based modelling strategies as shown in Table 7.
+Its design is centered on two architectural innovations that jointly enhance representational fidelity and computational efficiency:
+  1. 1.
+    Dual-branch temporal encoding.
+In contrast to recent CP-oriented architectures such as MCformer [10533212], which prioritise efficiency at the expense of temporal expressiveness, MSTN explicitly preserves multi-scale temporal structure.
+It deploys a parallel encoding strategy in which a global sequence-modeling pathway (Transformer or BiLSTM; CD strategy) captures long-range temporal dependencies, while a lightweight convolutional pathway (CI strategy; $\mathcal{O}(L)$) extracts fine-grained local temporal patterns.
+This duality directly addresses the multi-scale nature of real-world temporal signals, offering coverage beyond what is typically afforded by efficient CP-based methods.
+  2. 2.
+    Efficiency through ETA: collapsing $\mathcal{O}(L^{2})$ to $\mathcal{O}(1)$.
+MSTN applies the ETA mechanism immediately after the high-capacity encoders, collapsing the temporal dimension ($L\to 1$) via learned sequence aggregation.
+By performing the computationally intensive operations (e.g., $\mathcal{O}(L^{2})$ Transformer self-attention) before this aggregation step, the subsequent refinement layers—including feature fusion, squeeze–excitation recalibration, multi-head attention, and prediction modules—operate with a fixed $\mathcal{O}(1)$ cost with respect to $L$.
+This explicitly reshapes the model’s complexity profile, enabling MSTN to maintain the representational power of CD architectures while achieving substantially improved inference efficiency.
+4.3 Baselines
+This section presents the baselines used to assess MSTN on four time series tasks: (1) imputation, (2) long-term forecasting, (3) classification, and (4) generalizability study.
+4.3.1 Imputation
+Following the TimesNet [wu2023timesnettemporal2dvariationmodeling] protocol, we evaluate performance under the MCAR scenario with random masking ratios of $\{12.5\%,25\%,37.5\%,50\%\}$.
+Performance is quantified using MSE and MAE.
+The evaluation includes recent SOTA baselines such as GPT2(3) [zhou2023fitsallpowergeneraltime], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
 4.3.2 Long Term Forecasting
-The proposed MSTN model was evaluated on eight widely-used benchmark datasets: Weather, Traffic, Electricity, ILI, and four variants of ETT (ETTh1, ETTh2, ETTm1, ETTm2) [wu2022autoformer, wu2023timesnettemporal2dvariationmodeling, nie2023timeseriesworth64].
-These datasets are standard in the field and have been extensively used in prior studies, including EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime] and iTransformer [liu2024itransformerinvertedtransformerseffective].
-Notably, large-scale datasets such as Weather, Traffic, and Electricity contain a significantly higher number of time series, making their results more stable and less prone to overfitting compared to smaller datasets.
-We evaluate all methods across multiple prediction horizons $H\in{96,192,336,720}$ for standard benchmarks and $H\in{24,36,48,60}$ for the ILI dataset.
-4.2.2 Imputation
-The performance of imputation is assessed on benchmark datasets from the ETT [zhou2021informer], Electricity [electricityloaddiagrams20112014_321], and Weather [Kolle2025_BGCJenaWeather] domains, which commonly encounter missing data in real-world applications.
-Following the protocol established by TimesNet [wu2023timesnettemporal2dvariationmodeling], we evaluated performance under random masking ratios of $12.5\%,25\%,37.5\%,50\%$}.
-The evaluation includes recent SOTA baselines such as GPT2(3) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64], LightTS [zhang2022morefastmultivariatetime], and DLinear [zeng2022transformerseffectivetimeseries].
+The proposed MSTN model is evaluated for the long-term time series forecasting task.
+We evaluated all methods across multiple prediction horizons $H\in\{96,192,336,720\}$ for standard benchmarks and $H\in\{24,36,48,60\}$ for the ILI dataset.
+The forecast performance is quantified using MSE and MAE.
+The evaluation includes the following SOTA baselines: EMTSF [alharthi2025emtsfextraordinarymixturesotamodels], LLM4TS [10.1145/3719207], HiMTM [zhao2024himtmhierarchicalmultiscalemasked], TIME-LLM [jin2024timellmtimeseriesforecasting], MTST [zhang2024multiresolutiontimeseriestransformerlongterm], SOFTS [han2024softsefficientmultivariatetime], and iTransformer [liu2024itransformerinvertedtransformerseffective].
 4.3.3 Classification
-For classification tasks, we utilized ten multivariate benchmark datasets from the UEA time series classification archive [bagnall2018ueamultivariatetimeseries].
-These datasets encompass diverse applications including gesture and action recognition, heartbeat-based medical diagnosis, and other real-world scenarios.
+The performance of classification tasks is evaluated using the standard metric of classification accuracy.
 The evaluation compares with recent SOTA baselines such as GPT2(6) [alharthi2025emtsfextraordinarymixturesotamodels], TimesNet [wu2023timesnettemporal2dvariationmodeling], DLinear [zeng2022transformerseffectivetimeseries], ETSFormer [woo2022etsformer], FEDformer [zhou2022fedformer], and Informer [zhou2021informer].
-4.2.4 SOTA Baselines for Generalizability Study
-To rigorously evaluate the generalizability of the proposed MSTN framework, we conducted cross-domain benchmarking across seven publicly available international datasets.
-The evaluation encompasses datasets from diverse domains: human safety, with the Boubézoul et al.
-[BOUBEZOUL2020105577] (France) fall event dataset and the Rodegast et al.
-[DARUS33012023] (Germany) simulator for risky riding behaviour and collisions; general human activity, using UCI-HAR [har240] (Italy) and PAMAP2 [pamap2_physical_activity_monitoring_231] (USA) to assess robustness in physical activity recognition; animal behavior monitoring via the ActBe-Calf dataset [DISSANAYAKE2025111462] (Ireland) for welfare assessment; and mechanical systems, including MetroPT-3 [metropt3] (Portugal) for predictive maintenance and the NASA Turbofan Engine [NasaTurboFan] (USA) for sensor degradation forecasting.
-This selection tests model transferability across healthcare, activity recognition, agricultural technology, and industrial monitoring.
-All models were evaluated under a consistent five-fold cross-validation framework to ensure a robust and fair comparison, including recent architectures such as TimesNet [wu2023timesnettemporal2dvariationmodeling] and PatchTST [nie2023timeseriesworth64].
-4.3 Training and Model Configuration
-The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1+cu118 and trained on an NVIDIA T400 GPU (4 GB VRAM).
-The architecture processes variable-length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128→64 filters).
+4.3.4 Generalizability Study
+To evaluate generalizability between domains of the proposed MSTN framework, we benchmarked against baselines (e.g.
+TimesNet [wu2023timesnettemporal2dvariationmodeling], PatchTST [nie2023timeseriesworth64]) on seven publicly available international datasets spanning healthcare, activity recognition, agricultural technology, and industrial monitoring.
+All evaluations utilized a consistent five-fold cross-validation framework to ensure a robust and fair comparison.
+4.4 Experimental Setup
+4.4.1 Training and Model Configuration
+The proposed MSTN model is implemented in Python 3.13.1 using PyTorch 2.7.1 and trained on an NVIDIA T400 GPU (4 GB VRAM).
+The architecture processes variable length sequences through the sequence modeling cores BiLSTM (128 hidden units) or Transformer (4 layers, 8 attention heads), combined with a CNN pathway (128 $\to$ 64 filters).
 Training uses AdamW [kingma2017adammethodstochasticoptimization] optimizer with learning rate $3\times 10^{-4}$, batch size 64, and task-specific objectives (Focal Loss for classification, MSE for forecasting and imputation).
-Models train for up to 100 epochs with early stopping based on validation performance for task.
-Dataset-specific configurations include 96-step lookback for Traffic data and 50-step sequences for physical therapy datasets, with min-max feature normalization applied throughout.
+Models train for up to 100 epochs with early stopping based on validation performance.
+Data preprocessing employs min-max normalization to standardize feature distributions.
+For forecasting and imputation benchmarks, we utilize a fixed input lookback window of $L=96$ (except $L=36$ for ILI) across all prediction horizons, whereas for classification tasks involving variable length sequences (e.g., JapaneseVowels), inputs are standardized to the maximum sequence length (e.g., $L=29$) to facilitate consistent batch processing.
+4.4.2 Evaluation Metrics
+The performance of forecasting and imputation tasks is evaluated using two primary regression metrics that measure the discrepancy between predicted values and ground truth.
+For forecasting tasks, these metrics quantify the accuracy of future predictions, while for imputation tasks they assess the quality of missing value reconstruction.
+The performance is evaluated using:
+Mean Squared Error: $\quad\text{MSE}=\frac{1}{H}\sum_{i=1}^{H}(Y_{i}-\hat{Y}_{i})^{2}$ (15)
+Mean Absolute Error: $\quad\text{MAE}=\frac{1}{H}\sum_{i=1}^{H}|Y_{i}-\hat{Y}_{i}|$ (16)
+where $Y,\hat{Y}\in\mathbb{R}^{H\times C}$ are the ground truth and prediction results of the future with $H$ time points and $C$ channels.
+$Y_{i}$ denotes the $i$-th future time point.
+The performance of classification tasks is evaluated using four standard classification metrics that measure different aspects of predictive performance, particularly important for handling potential class imbalance in the datasets.
+The classification performance is evaluated using:
+$\text{Accuracy}=\frac{\text{TP}+\text{TN}}{\text{TP}+\text{TN}+\text{FP}+\text{FN}}$ (17)
+$\text{Precision}=\frac{\text{TP}}{\text{TP}+\text{FP}}$ (18)
+$\text{Recall}=\frac{\text{TP}}{\text{TP}+\text{FN}}$ (19)
+$\text{F1-Score}=2\times\frac{\text{Precision}\times\text{Recall}}{\text{Precision}+\text{Recall}}$ (20)
+where TP denotes True Positives, TN True Negatives, FP False Positives, and FN False Negatives.
 5 Acknowledgments
@@
 7 Author Contributions
-Sumit S Shevtekar led the research design, developed the MSTN architecture, implemented all software, conducted experiments across classification, forecasting and imputation tasks, performed data analysis and visualization, and wrote and edited the original draft.
-Chandresh K Maurya contributed to methodology design, experimental validation, and manuscript writing, review and editing.
-Gourab Sil participated in manuscript review and editing.
-All authors reviewed and approved the final manuscript.
-8 Competing Interests
-The authors declare that they have no known competing financial interests or personal relationships that could have appeared to influence the work reported in this paper.
+Sumit S Shevtekar led the research design, developed the architecture, implemented all software, conducted experiments, performed data analysis and visualization, and wrote and edited the original draft.
+Chandresh K Maurya contributed to experimental validation, manuscript writing, review, and editing